### v1

# Shuttling Compiler for Trapped-Ion Quantum Computers Based on Large Language Models

Abstract
Trapped-ion quantum computers based on segmented traps rely on shuttling operations to establish connectivity between multiple sub-registers within a quantum processing unit. Several architectures of increasing complexity have already been realized, including linear arrays, racetrack loops, and junction-based layouts. As hardware capabilities advance, the need arises for flexible software layers within the control stack to manage qubit routing—the process of dynamically reconfiguring qubit positions so that all qubits involved in a gate operation are co-located within the same segment. Existing approaches typically employ architecture-specific heuristics, which become impractical as system complexity grows. To address this challenge, we propose a layout-independent compilation strategy based on large language models (LLMs). Specifically, we fine-tune pretrained LLMs to generate the required shuttling operations. We evaluate this approach on both linear and branched one-dimensional architectures, demonstrating that it provides a foundation for developing LLM-based shuttling compilers for trapped-ion quantum computers.
1 Introduction
Shuttling-based trapped-ion quantum computers [KielpinskiMW02, KausahlLSHPSBMSP20, PinoDFGMABFHMRN21, MuraliDBM20, MosesB+23] are currently limited by both qubit count and the depth of reliably executable quantum circuits. Recent advances are gradually overcoming these barriers, enabling demonstrations of quantum advantage [AruteAB+19] and opening pathways toward practical applications in real-world computational tasks [PeruzzoMSYZLAO14, AnschuetzOAC19]. At the same time, this progress introduces additional hardware complexity, necessitating more sophisticated compilation strategies for efficient qubit routing and quantum circuit execution.
In trapped-ion systems, atomic ions are confined in segmented microchip traps [SchulzPSS06]. Each pair of electrodes generates a localized potential well that can hold multiple ions, effectively partitioning the trap into smaller, more manageable segments. The quantum information is encoded in the internal states of individual ions, each of which represents a qubit; accordingly, we use the term qubit to refer to an ion throughout the paper. By adjusting electrode voltages, trapping potentials can be dynamically reconfigured, enabling qubits to be shuttled between segments for successive gate operations. Segments are typically categorized as either storage segments, where qubits are held, or gate segments, where quantum operations are performed using laser or microwave radiation. In accordance with the current maturity level of trapped-ion hardware, the number of available gate segments is still very limited and often restricted to a single one [KausahlLSHPSBMSP20, SchmaleTBPKDOWB22].
Most current architectures arrange trap segments in a linear array [KausahlLSHPSBMSP20, PinoDFGMABFHMRN21, BowlerGLTHJHLW12, LeeJPJKC21] or in a racetrack loop [MosesB+23], where each segment connects to two neighbors. Such designs limit scalability, as the number of qubits scales linearly with the physical size of the trap. Introducing junctions [HensingerOSHYADMR06, BlakestadOVABLW09, WrightAFVDHPLDKSH13, ShuVBNVSB14] increases connectivity, allowing segments to connect to multiple neighbors. Quantinuum has demonstrated this idea in their Helios architecture, where a ring-shaped storage region is connected to two linear legs via an $\mathsf{X}$-shaped junction [Ransford+25]. However, Helios contains only a single junction. Increasing the number of junctions would enable more complex topologies in the future, such as two-dimensional grids [LekitschWFMDWH17].
Executing a quantum circuit on such hardware requires a sequence of physical operations. Because quantum gates can only be applied to qubits within the same gate segment, the trap must be dynamically reconfigured using four fundamental shuttling operations:
  • Translation: Moves the qubits from one trap segment to a neighboring segment.
  • Separation: Splits the qubits stored in one segment into two segments.
  • Merge: Combines qubits from two segments into a single segment.
  • Swap: Physically rotates the content of a segment by 180 degrees, reversing the order of the qubits within the segment.
While shuttling operations enable effective all-to-all connectivity between qubits and strongly suppress cross-talk errors, they also introduce substantial time overhead that lengthens operation timescales [HilderPOSOLRMSP22]. As a result, error sources such as dephasing and noise-induced heating become more pronounced, while the shuttling process itself can further degrade gate fidelities through ion heating. The gate fidelity quantifies how closely the implemented quantum gate matches its ideal operation, with lower fidelities indicating higher error probabilities during gate execution. Consequently, the number of shuttling operations must be minimized, highlighting the need for efficient compilation routines.
Compilation refers to the translation of a logical quantum circuit, which comprise abstract qubits and idealized gates, into a sequence of hardware-level instructions executable on a specific trapped-ion device. This process includes decomposing logical gates into the device’s native gate set, mapping logical to physical qubits, and determining a shuttling schedule that defines an executable sequence of shuttling operations. The resulting schedule must adhere to hardware constraints while minimizing resource overheads such as operation time, qubit transport distance, and cumulative error.
Existing shuttling compilers compute valid qubit routing schedules by generating shuttling schedules tailored to specific architectures, such as linear arrays [PinoDFGMABFHMRN21, Wagner22, DurandauWMBSPB23], racetrack loops [MosesB+23], drive-through architectures [ChangJCHL25], plus-sign layouts [SchmaleTBPKDOWB22], a ring with linear legs [Ransford+25], branched one-dimensional traps [KreppelMWHPSB24], and two-dimensional designs [WebberHWH20, DaiBR24, SchoenbergerHBW25, SchoenbergerHSW25, SchoenbergerW25]. However, as architectures grow more complex, designing separate heuristics for each layout becomes increasingly impractical.
To overcome the limitations of architecture-specific compilation, large language models (LLMs) offer a promising, layout-independent alternative. LLMs can perform new tasks from textual instructions or a few demonstrations [Brown+20], making them well suited for learning shuttling behavior directly from examples.
In this work, we investigate whether current LLMs can generate valid and efficient shuttling schedules and identify the challenges that arise when applying them to trapped-ion shuttling. We fine-tune LLaMA 3.2 [MetaAI24], Qwen 3 [QwenTeam25], DeepSeek LLM [Deepseek-AI24], and Gemma 3 [GemmaTeam25] on datasets derived from classical shuttling compilers for both linear [Wagner22, DurandauWMBSPB23] and branched one-dimensional architectures with junctions [KreppelMWHPSB24]. We then compare their performance across multiple benchmark circuits and additionally evaluate their ability to generalize to previously unseen trap layouts.
Our evaluation shows that the fine-tuned models were able to generate valid shuttling schedules and, in some cases, even surpassed classical shuttling compilers. However, performance degraded as the number of qubits and gates increased: all models frequently reached trap configurations that require numerous retries or prevent further schedule generation. Results on the unseen trap layouts further indicate that the models’ ability to generalize beyond the training architectures remains limited. Overall, the results show that our approach provides a solid foundation for developing LLM-based shuttling compilers.
Our compiler differs from machine-learning-based routing techniques developed for superconducting architectures [PalerSFA23, PozziHSM22, FanGL22, PascoalFA24, TangDKFKS24, SinhaAS22, RussoPPAC25, SangHH25, ZeynaliB25, CuomoCKTAPC23, EscofetOBPVFAAA25, SundaramGR25]. While those methods primarily optimize SWAP placement and gate parallelism on fixed coupling graphs, our approach accounts for the hardware-specific characteristics of trapped-ion devices. In particular, it considers shuttling operations, which are absent in superconducting systems, and the limited number of gate segments, which restrict the parallel execution of quantum gates. Our work is the first to employ large language models for compiling quantum circuits into shuttling schedules for trapped-ion quantum computers and to compare different LLMs for this task.
The remainder of this paper is structured as follows: Sec. 2 reviews prior work on qubit routing and computing shuttling operations. Sec. 3 presents our LLM-based approach, evaluated in Sec. 4. Finally, Sec. 5 summarizes our contributions and outlines future research directions.
2 Related Work
[…]
Within these compilation stacks, the individual compilation stages can be realized using a variety of algorithmic techniques. Compilation typically begins with a transformation to the hardware’s native gate set. Different methods pursue different optimization goals, such as reducing circuit depth, minimizing the number of specific gates, or adapting to particular architectures. Some compilers apply heuristics to obtain approximate solutions [YounisI22, JinHCHZZ23, KreppelMOWHPSB23, AroraXWLLDA25], while others employ SAT solvers to produce optimal results for smaller instances [PehamBKWB23]. More recently, machine learning has been applied to this task, ranging from recommender models [WeidenYKKI23] to reinforcement learning agents trained for specific optimization goals [MoroPRP21, FöselNML+22, Preti24, RietschDUPPMS24, Ruiz+25, RiuNVGE25, WeidenKKI25].
After this transformation, logical qubits must be mapped to physical qubits, and routing must be computed to enable the execution of gates under architectural constraints. This step is architecture-dependent, and techniques have been developed for a variety of platforms. For example, methods exist for neutral atom architectures [SchmidPW24, NottinghamPSWBCB24, StadeSVW24] and spin-qubit systems [ParaskevopoulosSAF23, ParaskevopoulosAF24], but most work focuses on superconducting and trapped-ion platforms.
In superconducting devices, architectures are modeled as coupling graphs, where vertices represent qubits and edges denote possible two-qubit interactions. Since two-qubit gates are restricted to adjacent qubits, additional SWAP gates are required to establish long-range connectivity. At the same time, gates on independent qubits can be executed simultaneously. Existing approaches differ in their optimization goals, such as minimizing SWAPs or maximizing gate parallelism. Proposed techniques span a wide range of methods, including heuristics and mathematical formulations [MaslovFM08, SaeediWD11, ShafaeiSP14, WilleLD14, LinSJ15, PedramS16, SiraichiSCP18, HattoriY18, LiDX19, MuraliBJCM19, ChildsSU19, CowtanDDKSS19, MatsuoY19, ZulehnerBW19, ZulehnerPW19, ItokoRIM20, NiuSST20, AlamAG20, TanC20, ZhouLF20, ZhangHQJYZ21, LiWSJDX22, ParkKKSK22, NanniciniBGJ22, HuaJCVKBLJZ23, PehamBW23, KhandavilliPNLND23, LiuYWHKI23, FuZWXSL23, ShaikP23, YangKSHD24, ChengYKWCH24, KattemölleH25], as well as machine-learning-based approaches. The latter include k-nearest neighbors, random forests, and multilayer perceptrons for qubit mapping [PalerSFA23], and reinforcement learning (RL) methods for routing [PozziHSM22, FanGL22, PascoalFA24, TangDKFKS24]. Specifically, [TangDKFKS24] employs a transformer-based neural network as an RL agent and reports a reduction in the number of SWAP gates by 10 to 20 % compared to the SABRE heuristic [LiDX19]. Hybrid approaches combining reinforcement learning with graph neural networks (GNNs) have also been proposed [SinhaAS22].
Beyond single-device architectures, several works address distributed superconducting architectures, where qubits are partitioned across multiple devices and minimizing inter-device qubit transfers becomes critical [RussoPPAC25, SangHH25, ZeynaliB25, CuomoCKTAPC23, EscofetOBPVFAAA25, SundaramGR25]. In this setting, GNN- and transformer-based models are used to capture global circuit structure [RussoPPAC25, SangHH25], and qubit reuse strategies are explored to further reduce communication overhead [SangHH25]. These approaches demonstrate reductions of up to 100 % in inter-device qubit transfers, effectively eliminating them when qubit reuse is enabled, and reductions of up to 99 % compared to QUBO [BandicPNORAVVAAF23] even without qubit reuse. The improvements are attributed to RL-based methods learning long-range dependencies across the entire circuit, enabling globally optimized qubit placements. Compared to Qiskit [WilleMN19], however, the reduction in inter-device transfers is in the range of 15 to 40 %, and some circuits still incur higher communication costs. Finally, [ZeynaliB25] proposes a noise-adaptive mapping strategy for distributed architectures that combines reinforcement learning for qubit placement with long short-term memory (LSTM) networks for noise migration and a multi-head self-attention mechanism to identify long-range qubit dependencies. Experimental results show that this approach increases the mean circuit fidelity by 49 %, raising it from 0.62 to 0.92, while reducing inter-chip operations by an average of 80 % compared to QUBO.
Trapped-ion quantum computers exhibit distinct hardware characteristics. In these systems, multiple qubits are confined together within individual trap segments. While quantum gates can only be applied at specific segments, the qubit distribution across segments can be dynamically reconfigured through shuttling operations. Shuttling compilers compute sequences of such operations, enabling circuit execution on a given trap. Starting from a quantum circuit, they map logical qubits to physical ones and then derive shuttling schedules. While the minimal number of timesteps can be computed using Boolean satisfiability methods [SchoenbergerHBW24], most existing compilers employ heuristics to reduce the number of shuttling operations. Prior work includes design studies analyzing how architectural choices affect shuttling overhead [MuraliDBM20, AshSakiTG22], as well as compilers targeting specific trap configurations as mentioned in Sec. 1. However, all these existing approaches are tightly coupled to specific architectures.
Beyond single-trap systems, compilation has also been explored for distributed trapped-ion architectures, where multiple traps are connected via photonic interconnects that mediate quantum communication between nodes [WuZWW25]. Such systems introduce additional scheduling and routing challenges, further highlighting the need for flexible compilation strategies that generalize across architectures.
In contrast to prior work, we present the first layout-independent shuttling compiler for trapped-ion quantum computers based on large language models (LLMs) and provide a comparison of different models for this task. By fine-tuning on examples of shuttling problems, LLMs are able to generalize beyond a single trap design and generate valid and efficient shuttling schedules for a variety of architectures, provided appropriate training data is available.
3 Compiler
This section explains how we fine-tune LLMs to generate shuttling schedules for quantum circuit execution. The process comprises three phases, illustrated in Fig. 1 and detailed in the following subsections.
Fig. 1: Overview of the three phases of our compiler. We first generate training and evaluation datasets in Alpaca format, where each entry consists of an instruction and its corresponding output. These datasets are then used to fine-tune an LLM with Axolotl, which involves pre-processing and training. Finally, the fine-tuned model is deployed for inference to compute shuttling schedules by submitting instructions to a vLLM server.
3.1 Dataset Generation
To fine-tune an LLM, we require examples that illustrate how shuttling is performed. These examples are provided as shuttling schedules, as shown in Fig. 2. Since the physical movements depend only on which qubits participate in each gate, rather than on the specific gate function, the gate functionality is omitted in this representation. Such schedules can be generated in different ways, for example by constructing random sequences, by execution guided through a reward function, or by employing another (not necessarily optimized) compiler. The number of required schedules depends on the trap architecture, its size, and the number of qubits. Larger and more complex traps, as well as higher qubit counts, demand more shuttling operations per gate. This increases both the difficulty for the LLM to learn generating valid shuttling schedules and the amount of data required for fine-tuning.
Fig. 2: Shuttling schedule of a quantum circuit with two qubits and three gates on a linear architecture consisting of three segments. Both qubits are initially located in gate segment 11. The lines indicate the qubit movements between the segments, while the purple dots denote gate executions. The executed shuttling operations are annotated alongside the lines. For dataset generation, the complete shuttling schedule is decomposed into multiple data entries, where each entry contains the shuttling operations between two consecutive gate executions, as indicated by the dotted lines.
The trap itself is modeled as a graph $G=(\mathcal{V},\mathcal{E})$, where the vertex set $\mathcal{V}$ represents the trap segments and the edge set $\mathcal{E}$ represents physical connections between neighboring segments. Each vertex is assigned a unique natural number for identification. In this work, we focus on linear architectures and branched one-dimensional architectures with junctions, for which example graphs are shown in Fig. 3.
(a) Graph representation of a linear architecture with 15 segments.
(b) Graph representation of a branched one-dimensional architecture with 23 segments. Junctions have a distance of three on each side of the gate segment, and each stack has a height of two.
Fig. 3: Graph representations of ion-trap architectures. Red vertices denote gate segments (GS), yellow vertices denote storage segments, and green vertices denote junctions, which are the only vertices connected to more than two others. Numbers serve as unique vertex identifiers.
From the generated shuttling schedules, we construct two datasets: one for training and one for evaluation. The evaluation dataset is smaller and is used during training to assess model performance on unseen data. Both datasets consist of multiple data entries, obtained by disassembling the shuttling schedules as illustrated in Fig. 2. Each entry corresponds to the portion of the schedule between two gate executions. The first entry of a schedule starts with the initial trap state and the full quantum circuit and contains the shuttling sequence up to the first gate execution. After this gate is executed, the trap enters a new state and the gate is removed from the circuit. The remaining operations then describe how this new state evolves toward executing the next gate. The second entry covers the operations until that execution, and so on, until all gates are removed and the entire schedule is decomposed into data entries. When constructing datasets, it is also possible to mix schedules from different architectures so that a single LLM can be fine-tuned across them.
Each data entry follows the Alpaca format [TaoriGZDLGLH23], consisting of an instruction and an output. Since the LLM can only process textual input, all required information is encoded as text.
3.1.1 Instruction
The instruction encodes all information the LLM requires to reason about the current trap configuration and valid operations. Since LLMs process purely textual input, all relevant information about the system must be encoded as text. For fine-tuning, the model input includes a graph-based description of the trap, its architectural constraints, the definitions of available shuttling operations, and the current trap state, including qubit locations, the circuit gates pending execution, and the shuttling operations permitted under that state.
The trap graph is defined by the sets of vertices and edges, where each vertex is identified by its unique natural number and each edge is represented by its incident vertices. Architectural constraints specify that each segment can hold a maximum number of qubits, which are arranged in a linear chain. Each qubit is assigned a position within this chain, allowing it to be uniquely referenced as a tuple $[v,p]$, where $v$ denotes the vertex number and $p$ the qubit’s position in the chain.
Fig. 4: Quantum circuit with four qubits and 14 gates. Purple gates form the first layer and are executable when their qubits are located in a gate segment, whereas yellow gates belong to deeper layers and are not yet executable. For each qubit, gates are numbered in ascending order to indicate their execution sequence.
In addition, the instruction specifies the quantum gates of the circuit, which must be executed on the given architecture. Each gate targets specific qubits and is assigned a unique natural number for identification. Due to architectural constraints, a gate is executable only if all involved qubits reside in the same gate segment, and this segment must not contain any additional qubits. As illustrated in Fig. 4, the gates are divided into two groups: gates in the first layer of the circuit which are executable when the corresponding qubits are in a gate segment and not yet executable gates in deeper layers. For each qubit, the gate numbers are assigned in ascending order. To ensure the gate execution in the correct order, gates targeting the same qubit must be executed in numerical order, while gates acting on different qubits may be executed in any order. Finally, to link the circuit qubits with the qubits in the trap, all qubits in the gate descriptions are replaced by tuples $[v,p]$ of their corresponding position in the trap.
To modify the trap state, the LLM requires a description of the available shuttling operations. As outlined in Sec. 1, these include the four shuttling operations Translate, Separate, Merge, and Swap, as well as gate execution. While gates can only be executed on the gate segments, depending on the hardware, separation, merging, and swapping may also be permitted on storage segments or a subset thereof.
In segmented trapped-ion architectures, junctions are formed by the intersections of electrodes, enabling qubits to be routed between different linear sections. Unlike storage segments, junctions do not provide a stable trapping potential suitable for holding qubits. Instead, the electrode potentials are shaped to guide qubits smoothly through the junction, ensuring efficient transport. Consequently, qubits must pass through junctions without stopping, and reversing direction within a junction is infeasible because the electric fields are engineered for unidirectional traversal rather than oscillatory motion [WrightAFVDHPLDKSH13, MokhberiSW17]. When computing shuttling operations, the LLM must respect this physical constraint. In particular, Separate, Merge, Swap, and gate execution cannot be performed on junction vertices, which must remain empty throughout execution. Furthermore, for Separate and Merge, the adjacent vertices involved must not be junctions. The operations are formally defined as follows:
  • Translate: Move all qubits from a vertex $v\in\mathcal{V}$ to a neighboring vertex $u\in\mathcal{V}$, provided that $\{v,u\}\in\mathcal{E}$ and $u$ is empty. After translation, $v$ becomes empty. To prevent qubits from being moved to and from a junction without first being moved to another storage segment, an additional rule applies: if $u$ is a junction, a translation from $u$ to $v$ cannot be reversed until a subsequent translation from $u$ to another neighbor $w\neq v$ has been executed.
  • Separate: Let $v\in\mathcal{V}$ be a vertex eligible for separation, containing $n\geq 2$ qubits, with empty neighbors v1v-1 and $v+1$. The qubits at positions $0,\dots,\lceil\tfrac{n}{2}\rceil$ in the qubit chain of $v$ are moved to v1v-1, while the remaining qubits are moved to $v+1$. After separation, $v$ is empty.
  • Merge: Let $v\in\mathcal{V}$ be an empty vertex eligible for merging, with neighbors v1v-1 and $v+1$ containing $m$ and $n$ qubits, respectively, such that $m+n$ does not exceed the maximum segment capacity. The qubits from v1v-1 and $v+1$ are moved into $v$, where positions $0,\dots,m-1$ are occupied by the qubits from v1v-1 and positions $m,\dots,m+n-1$ by the qubits from $v+1$. After merging, v1v-1 and $v+1$ are empty.
  • Swap: Let $v\in\mathcal{V}$ be a vertex eligible for swapping, containing at least two qubits. Reverse the order of the qubit chain in $v$.
  • Execute Gate: Perform one of the first-layer gates in a gate segment, provided all previously defined execution conditions are satisfied.
After specifying all rules to the LLM, the overall goal is defined: execute all gates using the minimal number of operations. To accomplish this, the LLM must also be aware of the current trap state. Accordingly, the instruction includes four enumerations:
  1. 1)
    Qubit positions: All qubits in the trap, each represented by its position tuple $[v,p]$ as described above.
  2. 2)
    First-layer gates: All gates in the first layer of the circuit that are executable when their corresponding qubits reside in a gate segment, along with the tuples of the qubits they act on.
  3. 3)
    Deeper-layer gates: Gates in deeper layers that are not yet executable, along with the tuples of the qubits they act on. To keep the focus on near-term execution, this list includes only the gates that become executable next, rather than all remaining gates in the circuit.
  4. 4)
    Allowed shuttling operations: All shuttling operations currently valid for the given trap state, allowing the LLM to select a valid first step without ambiguity.
The instruction concludes by defining the task for the LLM: generate a shuttling sequence that enables the execution of one of the first-layer gates, contributing to the overall goal. After each operation, the LLM should also provide the updated qubit positions, the list of first-layer gates with updated position tuples of the qubits, and the new set of allowed shuttling operations. Requiring these additional outputs at each step supports the LLM to reason carefully about the current state and the implications of each operation.
3.1.2 Output
The task of the LLM is to generate a shuttling sequence which allows to execute one of the first-layer gates. To learn the generation of such sequences, the LLM needs examples of expected shuttling sequences during the training phase that solve the task given in the instruction. Therefore the output contains the part of the previous generated shuttling schedule that shows how one of the first-layer gates can be executed from the given state. As long as not all requirements are fulfilled to execute one of the executable gates, usually multiple shuttling operations are required in order to execute one of these gates. The number of operations depend on the size of the trap and the number of qubits used. The reason is that with increasing trap size and qubit number, more shuttling operations are required to reconfigure the trap segments so that the trap is in a valid state for a gate execution.
The output begins with the first shuttling operation which is to be executed based on the state given in the instruction. Since the instruction demands also to output the updated qubit positions and based on this the first-layer gates and the new allowed shuttling operations after each operation, these information is also contained and presented in the same way as in the instruction. This information helps the LLM to understand which changes happen in the trap state based on the executed shuttling operation and which consequences they have. Afterwards, the output contains the next shuttling operation based on the new state. This continues until the operation “Execute Gate” indicates that one of the first-layer gates is to be executed.
After generating the training and evaluation datasets, we use them to fine-tune an LLM with Axolotl [Axolotl23].
3.2 Fine-tuning with Axolotl
After dataset preparation, we fine-tune pre-trained LLMs from Huggingface [Wolf+20] to compute shuttling schedules. For this, we employ Axolotl [Axolotl23], which provides seamless compatibility with Huggingface models. Axolotl supports various training approaches, including full fine-tuning, Low-Rank Adaptation (LoRA) [HuSWALWWC22], and Quantized LoRA (QLoRA) [DettmersPHZ23]. Since LoRA and QLoRA did not yield the expected results (see Sec. 3.4), we performed full fine-tuning of the LLM. Furthermore, Axolotl offers performance optimizations such as multipacking [KunduLWGM24], FlashAttention 2 [Dao24], sequence (context) parallelism [NarayananSCLPKVKBCPZ21], and multi-GPU training with DeepSpeed [RasleyRRH20], which we describe in more detail below.
3.2.1 Performance Optimizations
To minimize resource requirements and enable fine-tuning on as few GPUs as possible, it is essential to reduce memory usage and parallelize execution. We achieve this with DeepSpeed, employing the Zero Redundancy Optimizer (ZeRO) at optimization level 3 [RajbhandariRRH20] to accelerate and distribute fine-tuning [ShoeybiPPLCC19]. In this setting, 16-bit model parameters are partitioned across GPUs, while DeepSpeed manages parameter partitioning and collection during forward and backward passes. To further reduce memory consumption, we use bfloat16 as the 16-bit floating-point format. Additionally, we apply ZeRO Offloading [RenRARYZLH21], which transfers optimizer states, as well as model parameters and gradients when inactive, to the host CPU and RAM.
The data entries in our datasets vary considerably in length due to differing output sizes, which depend on the number of shuttling operations required to execute one of the first-layer gates. While in some cases, a gate can be executed directly within the current trap configuration, other cases require complex shuttling operations, producing significantly longer outputs. Since fine-tuning requires sequences of uniform length, shorter sequences are padded with special tokens. To avoid excessive padding to match the longest sequence, we apply multipacking, which enables training on multiple shorter entries simultaneously, provided the combined length does not exceed the model’s context window. To efficiently process longer sequences, we employ FlashAttention 2, an optimized attention mechanism that improves memory efficiency and accelerates training, thereby allowing for longer sequence lengths. For cases where individual sequences exceed the VRAM capacity of a single GPU, we use sequence parallelization. This approach relies on blockwise ring attention [LiuZA24], which enables distributed computation of self-attention and feedforward layers across multiple GPUs, while overlapping communication and computation to minimize overhead.
Moreover, because each instruction in our dataset specifies a problem and its context, while the corresponding output provides a step-by-step solution, we compute the loss exclusively on the output. This encourages the model to generalize solutions rather than memorize instructions.
The fine-tuning process is divided into two phases: pre-processing and training. These phases are managed automatically by Axolotl upon submission of a YAML configuration file, which specifies, among other parameters, the performance optimization settings and the file paths to the model, training and evaluation datasets, and the DeepSpeed configuration. The latter defines all DeepSpeed-related settings.
In the following, we discuss the two fine-tuning phases in more detail.
3.2.2 Pre-processing
The pre-processing phase prepares the dataset to streamline subsequent training. During this phase, Axolotl validates the YAML configuration file, ensuring that all required libraries are installed and that the selected optimizations are compatible. It then initializes the tokenizer and parses and formats both the training and evaluation datasets. Since the loss is computed exclusively on the output, instruction tokens are masked. With multipacking, multiple data entries are concatenated up to a defined maximum length with respect to the correct loss mask. The processed dataset, including tokenized inputs, loss masks, and associated metadata, is saved in Apache Arrow format for use during training.
3.2.3 Training
During the training phase, Axolotl coordinates the various optimization strategies. It loads the model, tokenizer, and pre-processed datasets using Huggingface’s datasets library and distributes them according to the selected optimizations. When loading the model, Axolotl integrates the memory-efficient attention mechanism from Flash Attention 2. It also configures Huggingface’s Trainer from its transformer library with the required parameters and connects it to DeepSpeed. While DeepSpeed handles memory optimizations and parallelization, the Trainer manages the training loop, which loads micro-batches of tokenized data. To conserve memory, each iteration uses a batch size of one; the forward pass computes the loss on the output, which is backpropagated to calculate gradients. We use gradient accumulation over eight batches to mitigate gradient noise and to improve convergence despite the small batch size. The accumulated gradients are applied by the AdamW optimizer [LoshchilovH19] to update the network weights. In addition to training, the Trainer handles logging, evaluation at regular intervals, and checkpointing. We fine-tune the model for a single epoch on the training dataset. Performing a single epoch rather than multiple ones enables us to leverage larger datasets, so the model encounters a greater variety of examples. Upon completion, the fine-tuned model is saved for inference.
3.3 Inference with vLLM
After fine-tuning, the model is deployed for inference using vLLM [KwonLZSZYGZS23], which provides efficient multi-GPU support. Its PagedAttention mechanism optimizes memory management of attention keys and values, thereby reducing overhead. Since vLLM is fully compatible with Huggingface models, it integrates seamlessly with our fine-tuned model. For inference, we launch a local vLLM server, which automatically loads the model and its tokenizer. The server exposes an OpenAI-compatible API accepting JSON-formatted HTTP requests. Each request contains an instruction in the same format as the training data, together with the maximum number of tokens to generate. Since outputs can be lengthy depending on the required operations, we set this limit to 29,000 tokens. The server returns the response in JSON format, from which we extract the generated output. To obtain clean text, special tags such as <think> are removed.
Each request to the vLLM server produces a single output describing the shuttling operations required to execute one of the first-layer gates. Thus, multiple requests are necessary to construct a complete shuttling schedule. Since outputs may include invalid operations inconsistent with the trap architecture or current state, we developed a Python script to automate schedule generation and validation. The script tracks both the qubit positions in the trap and the remaining, not-executed gates of the circuit. At initialization, it loads the quantum circuit, heuristically determines an initial qubit placement on the trap graph, and identifies first-layer gates, deeper-layer gates, and allowed shuttling operations. Using this information, it constructs an instruction in the same format as the training data and submits it to the vLLM server.
The returned output is then validated by checking whether all operations are consistent with the architectural rules and whether at least one first-layer gate has been successfully executed. If the output is valid, redundant translations, such as unnecessary back-and-forth movements of qubits between segments, are removed to reduce the overall operation count. Similarly, consecutive merges and separations on the same qubits without intermediate operations, as well as successive swap operations, are eliminated. Afterwards, the script updates the qubit positions, removes the executed gate from the circuit, and generates the next instruction based on the updated state before resubmitting it to the vLLM server. This iterative process continues until all gates have been executed. If an invalid output is produced, the same instruction is resubmitted to the vLLM server to obtain a valid result. To prevent infinite loops, the procedure terminates if ten consecutive invalid outputs occur for the same instruction.
In principle, it would also be possible to request multiple outputs from the vLLM server at once, validate and post-process them, and then select the valid output with the fewest operations. However, we chose to generate and evaluate only one output at a time. The reasons are twofold: first, depending on the complexity of the shuttling operations required to perform one of the first-layer gates, generating each output may take over one minute; and second, always selecting the locally optimal output with the fewest operations does not guarantee a globally minimal shuttling schedule. The process may converge to a local minimum, requiring additional operations to escape. Avoiding this would require exploring multiple valid continuations for each output over several steps and then selecting the output that starts the path with the lowest total number of operations. While such a lookahead strategy could reduce the overall number of operations, its runtime would grow exponentially with the number of steps.
3.4 Alternative Approaches and Challenges
In addition to the fine-tuning strategy described above, we explored several alternative approaches to reduce data complexity and resource requirements. While these methods provided insights, they did not yield satisfactory results in practice.
First, to reduce the length of the data entries, we attempted to reduce the output size in the training and evaluation datasets by splitting each shuttling operation into a separate entry. In this setup, the first entry contained the same state information as the original instruction but produced only the first shuttling operation as output. After applying this operation, the qubit positions were updated, and the new state was used in the instruction of the next entry, whose output contained the second shuttling operation. This is continued until all operations of the original entry were separated. For the example in Fig. 2, this would result in nine entries instead of three. While this approach reduced the size of individual entries, it substantially increased their total number. Moreover, during inference with models fine-tuned on such datasets, the LLM often selected randomly among the valid shuttling operations. Although the chosen operations were valid, they did not necessarily advance the execution of the quantum circuit.
Second, instead of limiting the instruction to the first-layer gates and a subset of deeper-layer gates, we included the entire remaining circuit in OpenQASM format [CrossBSG17, CrossJADBHRSSGJ22], along with a mapping from qubits to their positions in the trap. While this approach has much longer state descriptions, it also introduced a critical drawback: the LLM was unable to reliably identify the first-layer gates and frequently generated shuttling operations that targeted arbitrary, and often deeper-layer, gates.
Finally, we evaluated parameter-efficient fine-tuning methods such as LoRA and QLoRA, which restrict training to the linear layers and thereby reduce the number of trainable parameters and VRAM requirements. However, models fine-tuned with LoRA or QLoRA tended to generate step-by-step reasoning instead of directly producing the expected output format. These derivations often failed to fully account for the rules governing shuttling operations, leading to incorrect schedules. In addition, the deviation from the expected output format made it more difficult to automatically extract the relevant shuttling operations.
Given these limitations, we focused on the main fine-tuning approach described above.
4 Evaluation
This section evaluates the performance and generalization capabilities of our LLM-based shuttling compiler. We fine-tuned several LLMs using shuttling examples from both the linear architecture and the branched one-dimensional architecture with junctions, and then assessed how well the resulting models generate valid and efficient shuttling schedules. The evaluation includes multiple benchmark circuits compiled for the two architectures seen during training as well as three previously unseen layouts, shown in Fig. 5. The following subsections describe the experimental setup and present the results.
Fig. 5: Graph representations of ion-trap architectures used to evaluate the generalization capabilities of the fine-tuned LLMs on previously unseen layouts. Architectures (a) and (b) feature three-way junctions, while architecture (c) contains four-way junctions. For each layout, the number of yellow storage vertices was scaled to match the qubit count of the respective test circuit.
4.1 Setup
In the following, we first describe the experimental setup used for fine-tuning the models and then outline the setup employed for inference with the fine-tuned LLMs.
4.1.1 Fine-tuning
We performed fine-tuning on a single node equipped with eight H100 GPUs, each providing 80 GB of VRAM. Therefore, first we generated separate training and evaluation datasets containing examples for both the linear and the branched one-dimensional architectures with junctions. In both architectures, a single gate segment was used, and the separate, merge, and swap operations were executable only within this segment. This configuration imposes the strongest execution constraints, making it the most challenging setup for the LLMs to learn. The trap graphs were designed such that the number of storage segments on each side of the gate segment matched the number of qubits in the circuit.
For the linear architecture, the number of qubits was varied between 2 and 16. For the junction architecture, the number of qubits ranged from 3 to 16, and we additionally varied the stack depth, specifying the number of storage segments per stack, and the junction distance, defining the spacing between the junctions on each side of the gate segment, as illustrated in 3(b). In the following, we denote the stack depth by $s$ and the junction distance by $d$. Due to hardware limitations, we restricted the datasets to circuits of these sizes, as larger circuits would require more complex shuttling patterns and, consequently, additional VRAM and GPUs for fine-tuning.
Shuttling schedules were generated using the compilers from [Wagner22] for the linear architecture and [KreppelMWHPSB24] for the branched architecture. The underlying quantum circuits were randomly generated with Qiskit [WilleMN19], containing single- and two-qubit gates with a depth of five. These circuits were then compiled into a trapped-ion-compatible format using the compiler from [KreppelMOWHPSB23].
For the linear architecture, 120 shuttling schedules per qubit count were included in the training dataset and 30 in the evaluation dataset. For the junction architecture, eight schedules per trap design and qubit count were used for training and two for evaluation. Both datasets contained approximately equal proportions of examples from each architecture and were automatically shuffled by Axolotl prior to fine-tuning to ensure uniform data distribution.
Due to the hardware constraints and the size of the dataset entries, model sizes were limited to 50 GB, preventing us from fine-tuning large models such as DeepSeek V3 [Deepseek-AI24_2], DeepSeek R1 [GuoYZ+25], or LLaMA 4 [MetaAI25]. Within these limits, we fine-tuned a variety of models, including LLaMA 3.2 with 3 billion parameters [MetaAI24], Qwen 3 with 4 and 14 billion parameters [QwenTeam25], DeepSeek LLM with 7 billion parameters [Deepseek-AI24], and Gemma 3 with 27 billion parameters [GemmaTeam25].
In addition, we attempted to fine-tune GPT-OSS with 20 billion parameters [OpenAI25], which required replacing DeepSpeed with FSDP2 [Zhao+23]. However, we were unable to evaluate this model on the test set because it consistently produced invalid responses after fine-tuning: the first one or two shuttling operations were always truncated, making it impossible to construct a valid schedule.
4.1.2 Inference
For inference, we evaluated LLaMA, Qwen, and DeepSeek on a single A100 GPU with 40 GB VRAM on the Mogon NHR cluster of Johannes Gutenberg University. Due to its larger model size, Gemma was evaluated on a single A100 GPU with 80 GB VRAM on the Mogon KI cluster. As test circuits for compilation with the finetuned LLMs, we selected three of the 153 circuits from the library of [Xiangzhen19], which have been used in prior work [ZhouLF20, CowtanDDKSS19, KreppelMOWHPSB23, KreppelMWHPSB24]: 4mod5-bdd_287 (7 qubits, 106 gates), mini_alu_305 (10 qubits, 263 gates), and cnt3-5_179 (16 qubits, 230 gates). The gate counts refer to the circuits after compilation using [KreppelMOWHPSB23]. We chose these circuits because they span different qubit counts and, within the library, have the lowest gate counts for their respective qubit numbers. Since each request to the vLLM server returns the shuttling operations for a single gate, inference time is strongly influenced by the number of gates. Thus, selecting circuits with minimal gate counts kept the benchmarks manageable while still allowing us to evaluate performance across multiple qubit regimes.
For each circuit, we used the fine-tuned models to generate shuttling schedules for the linear architecture and for 10 (4mod5-bdd_287), 18 (mini_alu_305), and 29 (cnt3-5_179) branched architectures, each defined by different combinations of stack depth and junction distance. To assess generalization, we additionally evaluated the models on the previously unseen architectures shown in Fig. 5. As in the training data, each graph contained a single gate segment, and the separate, merge, and swap operations were executable only within that segment. We also evaluated the stochastic behavior of the models by varying the temperature $T$, which controls how deterministic the output is. Lower temperatures yield more deterministic outputs, with $T=0.0$ making the model behave fully deterministically. In contrast, higher temperatures increase the level of randomness in the generated outputs. To balance determinism and variability, we evaluated each model–circuit–architecture combination at $T=0.7$ and $T=1.0$. For each configuration, we performed ten runs and limited the runtime to eight hours per run.
Tab. 1: Operation counts of shuttling schedules generated by LLaMA, Qwen, DeepSeek, and Gemma with different temperatures $T$ compared to the classical compiler. For each model, only the run with the fewest operations is reported. For the junction architecture, we list only the best-performing configuration and specify the corresponding stack depth $s$ and junction distance $d$ as $(s,d)$. If none of the ten runs succeeded, the entry is marked as “failed”, with the number of successfully executed gates shown in parentheses. If all junction configurations failed, the table reports the configuration that completed the most gates before failure.
Archi- tecture Classical LLaMA 3.2-3B Qwen 3-4B Qwen 3-14B DeepSeek-7B Gemma 3-27B
$T=0.7$ $T=1.0$ $T=0.7$ $T=1.0$ $T=0.7$ $T=1.0$ $T=0.7$ $T=1.0$ $T=0.7$ $T=1.0$
Circuit: 4mod5-bdd_287, 7 qubits, 106 gates
Linear 607 559 653 1,071 517 842 701 598 543 700 825
Best 482 506 464 547 537 518 545 541 566 533 521
junction $(1,3)$ $(2,1)$ $(2,3)$ $(2,2)$ $(1,2)$ $(2,2)$ $(2,2)$ $(3,1)$ $(2,1)$ $(2,3)$ $(2,2)$
5(c) failed 569 failed failed failed failed failed failed failed failed
(12) (23) (33) (52) (26) (4) (4) (7) (16)
Circuit: mini_alu_305, 10 qubits, 263 gates
Linear 2,390 failed failed failed 3,124 failed failed failed failed failed failed
(106) (161) (165) (95) (86) (123) (98) (70) (75)
Best junction 1,624 1,791 1,809 failed 2,776 failed failed failed failed failed failed
(168) (85) (70) (67) (69) (95) (126)
$(2,1)$ $(2,3)$ $(2,2)$ $(4,2)$ $(5,1)$ $(3,2)$ $(1,3)$ $(4,2)$ $(2,4)$ $(4,2)$ $(5,1)$
5(c) failed failed failed failed failed failed failed failed failed failed
(38) (48) (19) (13) (15) (15) (0) (0) (16) (13)
Circuit: cnt3-5_179, 16 qubits, 230 gates
Linear 3,467 failed failed failed failed failed failed failed failed failed failed
(142) (146) (103) (118) (80) (98) (63) (59) (88) (43)
Best junction 1,716 2,435 2,017 failed failed failed failed failed failed failed failed
(140) (111) (64) (59) (44) (49) (56) (53)
$(2,1)$ $(1,1)$ $(3,1)$ $(10,1)$ $(3,2)$ $(6,3)$ $(6,1)$ $(4,3)$ $(5,4)$ $(4,3)$ $(4,4)$
5(c) failed failed failed failed failed failed failed failed failed failed
(12) (16) (7) (9) (4) (3) (5) (5) (5) (7)
4.2 Results
This subsection presents the results of our evaluation. The minimal number of shuttling operations required by the schedules generated by LLaMA, Qwen-4B, Qwen-14B, DeepSeek, and Gemma is summarized in Tab. 1, together with the corresponding values from the classical compiler. If a model failed to produce a valid schedule, the table reports how many gates were executed before the generation process became invalid. Since schedule generation on the architectures shown in 5(a) and 5(b) failed at early stages, these architectures are not included in the table.
Additional statistics on the generation process are provided in Tab. 2, including the number of required retries and the number of response tokens. We measure the number of generated tokens rather than runtime because token counts are hardware-independent and commonly used as a billing metric for LLMs [VelascoTOG2025]. In contrast, runtime depends heavily on hardware factors such as GPU type, the number of GPUs used, their distribution across cluster nodes, and the intra- and inter-node interconnects. For each configuration, we report both the token count of the final valid schedule, which excludes retries, and the total tokens generated across all attempts, including failed ones. The table also indicates, for each junction architecture, the number of configurations that produced valid schedules and how many of these schedules required fewer shuttling operations than the classical compiler with the same configuration.
In the following, we first present the results obtained by the LLMs for the linear and branched one-dimensional architectures with junctions, comparing them with each other and with the classical compiler. Subsequently, in Sec. 4.2.2, we evaluate the performance of the LLMs on the previously unseen architectures depicted in Fig. 5.
Tab. 2: Statistics for the best runs reported in Tab. 1. For each run, we report the number of retries and the number of response tokens required to generate the complete shuttling schedule. The token count is given once without retries, where only the tokens of the final valid solution are included, and once with retries, where all tokens generated during invalid attempts are also counted. For the junction architecture, we additionally report how many configurations produced valid schedules and, among those, how many achieved fewer operations than the classical compiler on the same configuration.
Archi- tecture Parameter LLaMA 3.2-3B Qwen 3-4B Qwen 3-14B DeepSeek-7B Gemma 3-27B
$T=0.7$ $T=1.0$ $T=0.7$ $T=1.0$ $T=0.7$ $T=1.0$ $T=0.7$ $T=1.0$ $T=0.7$ $T=1.0$
Circuit: 4mod5-bdd_287
Retries 9 9 19 26 2 17 5 15 17 15
Resp. toks 61,026 78,302 143,057 60,440 99,779 85,658 75,542 72,603 85,898 103,769
Linear w/o retries
Resp. toks 183,015 282,865 465,466 372,417 179,779 416,445 102,256 172,337 250,340 271,816
w/ retries
Retries 3 1 17 15 22 14 28 40 9 25
Best junction Resp. toks 65,514 56,347 79,480 78,681 98,207 80,583 90,623 93,918 76,700 73,839
w/o retries
Resp. toks 74,007 57,120 238,377 226,940 364,637 176,697 463,023 389,144 169,306 239,441
w/ retries
Succ. junction confs. 10/10 10/10 8/10 8/10 9/10 9/10 1/10 5/10 5/10 7/10
Confs. outperf. class. 0/10 3/10 0/8 1/8 1/9 1/9 0/1 0/5 0/5 0/7
Retries 36
Resp. toks 67,745
5(c) w/o retries
Resp. toks 228,928
w/ retries
Circuit: mini_alu_305
Retries 33
Resp. toks 555,430
Linear w/o retries
Resp. toks 1,251,636
w/o retries
Retries 26 29 58
Best junction Resp. toks 401,431 392,065 629,391
w/o retries
Resp. toks 725,633 766,859 2,187,631
w/ retries
Succ. junction confs. 15/18 11/18 0/18 1/18 0/18 0/18 0/18 0/18 0/18 0/18
Confs. outperf. class. 1/15 2/11 0/1
Circuit: cnt3-5_179
Retries 27 104
Best junction Resp. toks 646,571 564,133
w/o retries
Resp. toks 1,415,626 2,592,367
w/ retries
Succ. junction confs. 1/29 2/29 0/29 0/29 0/29 0/29 0/29 0/29 0/29 0/29
Confs. outperf. class. 0/1 0/2
4.2.1 Linear and branched one-dimensional architectures
For 4mod5-bdd_287 on the linear architecture, DeepSeek, LLaMA at $T=0.7$, and Qwen-4B at $T=1.0$ produced schedules requiring fewer operations than the classical compiler, whereas Qwen-14B and Gemma consistently generated schedules with higher overhead. The best result was obtained by Qwen-4B at $T=1.0$, reducing the operation count by 17 %, indicating that the smaller Qwen model performs better than the larger one on this architecture. However, Qwen-4B also yielded the worst result at $T=0.7$, requiring 76 % more operations than the classical baseline. Across all models, 2 to 26 retries were needed, and valid schedules contained between 60,440 and 143,057 response tokens. When retries were included, token counts increased by up to a factor of 6.2. Notably, the best-performing Qwen-4B run also produced the smallest final token count, but its high number of 26 retries led to the largest overall increase when considering all generated tokens.
For the junction architecture, LLaMA produced valid schedules for all ten tested configurations and, at $T=1.0$, outperformed the classical compiler in three cases. Qwen-14B produced valid schedules for all configurations except $s=1$, $d=4$ and surpassed the classical baseline for $s=5$, $d=1$ at both temperatures. Qwen-4B generated valid schedules for eight configurations and outperformed the classical compiler for the same configuration as Qwen-14B, but only at $T=1.0$. Gemma yielded valid schedules for seven configurations at $T=1.0$ and five at $T=0.7$, while DeepSeek produced valid schedules for only five configurations at $T=1.0$ and one at $T=0.7$; neither model surpassed the classical baseline. LLaMA’s best result, obtained for $s=2$, $d=3$ at $T=1.0$, required 4 % fewer operations than the best classical configuration $s=1$, $d=3$ and required only a single retry, leading to similar token counts with and without retries. In contrast, the best results of the other models required 7 % to 12 % more operations than the classical compiler and involved substantially more retries, ranging from 9 to 40. This resulted in token counts up to 5.1 times higher when retries were included.
As the number of qubits increases, generating valid schedules becomes substantially more challenging. While LLaMA was still able to produce valid schedules for several junction configurations on both larger benchmark circuits, Qwen-4B generated a valid schedule only for the linear architecture and a single junction configuration of mini_alu_305. However, Qwen-14B, DeepSeek, and Gemma failed to generate any complete schedules for the two larger circuits.
For mini_alu_305 on the linear architecture, Qwen-4B at $T=1.0$ was the only model to produce a complete schedule, requiring 3,124 shuttling operations, corresponding to a 31 % increase over the classical compiler. The resulting schedule contained 555,430 tokens, and its computation required 33 retries, increasing the total number of generated tokens by a factor of 7.7. All other models, including Qwen-4B at $T=0.7$, failed to generate a complete schedule for this circuit. The increasing difficulty can be attributed to the larger number of qubits, which makes trap reconfiguration more complex and often necessitates recursive shuttling operations. Consequently, the models frequently reach trap configurations from which continuing schedule generation becomes infeasible. Moreover, since our compiler generates schedules gate by gate, the likelihood of encountering unrecoverable configurations grows with circuit depth. Nevertheless, the models were able to execute between 70 and 165 of the 263 gates, indicating that substantial partial progress remains achievable. Among these models, Gemma performed worst, whereas LLaMA executed more than 100 gates before failing.
On the junction architecture, Gemma, Qwen-14B, and DeepSeek again failed to complete any schedule, executing at most 126, 85 and 69 gates, respectively. Qwen-4B produced a valid result only for the configuration $s=5$, $d=1$ at $T=1.0$, requiring 71 % more operations than the best classical configuration $s=2$, $d=1$ with 1,624 operations. This run required 58 retries, increasing the total number of generated tokens to 3.5 times the size of the final schedule. Among the failed runs, Qwen-4B at $T=0.7$ performed best, executing 168 gates before failure for the configuration $s=4$, $d=2$. However, LLaMA performed substantially better: of the 18 tested configurations, 15 at $T=0.7$ and 11 at $T=1.0$ yielded valid schedules, with one and two of these, respectively, outperforming the classical compiler. Unlike the linear architecture, where LLaMA failed, the presence of stacks in the junction architecture provides intermediate storage locations that simplify reconfiguration and reduce the need for recursive operations. Relative to the best classical configuration, LLaMA achieved the best results for the configurations $s=2$, $d=3$ at $T=0.7$ and $s=2$, $d=2$ at $T=1.0$ requiring only 11 % more operations. These runs required 26 and 29 retries, respectively, roughly doubling the number of generated tokens compared to the final schedules.
For cnt3-5_179, none of the models generated a complete schedule on the linear architecture. LLaMA executed more than 140 gates before failing, whereas Qwen-4B failed after over 100 gates and Qwen-14B completed up to 98 gates. DeepSeek failed after approximately 60 gates. Gemma’s performance strongly depended on the temperature: at $T=0.7$ it failed after executing 88 gates, while at $T=1.0$ it executed only 43 gates before failing. Notably, LLaMA’s performance indicates that, in principle, shuttling schedules for 16-qubit circuits are attainable, although robustness remains a limiting factor.
On the junction architecture, both Qwen models, Gemma and DeepSeek failed for all configurations. While Qwen-4B could execute up to 140 gates, only about 60 gates for Qwen-14B, 55 for Gemma, and 45 for DeepSeek could be executed. As for mini_alu_305, LLaMA also performed best here. Out of 29 tested configurations, LLaMA produced valid schedules for one at $T=0.7$ and two at $T=1.0$, but none outperformed the classical compiler. The best classical configuration $s=2$, $d=1$ required 1,716 operations, whereas LLaMA’s best result achieved with configuration $s=3$, $d=1$ at $T=1.0$ required 18 % more operations. This run required 104 retries, increasing the number of generated tokens by a factor of 4.6 relative to the final schedule.
4.2.2 Unseen architectures
In addition to linear and branched one-dimensional architectures with junctions, we evaluated the generalization capabilities of the fine-tuned LLMs on the unseen architectures shown in Fig. 5. While the layouts in 5(a) and 5(b) feature three-way junctions, similar to the branched architectures used during training, the architecture in 5(c) includes four-way junctions.
For the architectures shown in 5(a) and 5(b), none of the models generated a valid shuttling schedule for any of the evaluated circuits. While on the first architecture, all models failed immediately during schedule generation, on the second one, the models were able to execute up to four gates before failing.
In contrast, one model succeeded in generating a valid schedule for the architecture shown in 5(c) for the circuit 4mod5-bdd_287. LLaMA at $T=0.7$ produced a valid schedule with 569 shuttling operations but required 55 retries, increasing the total token count by a factor of 6.0 relative to the final schedule. At $T=1.0$, the model executed 74 of the 106 gates before failing. All other models failed to generate valid schedules on this architecture. Nevertheless, partial progress was observed: Qwen-14B executed up to 52 gates, Qwen-4B up to 33 gates, and Gemma at $T=1.0$ 16 gates, while Gemma at $T=0.7$ and DeepSeek failed at an early stage.
Consistent with the results on linear and branched architectures, performance on the four-way junction architecture degraded as the number of qubits increased. For the circuit mini_alu_305, no model generated a valid schedule. LLaMA executed up to 48 gates before failing, whereas both Qwen models and Gemma executed approximately 15 gates. DeepSeek performed worst, failing to execute even a single gate. Excluding DeepSeek, these results indicate that limited progress remains possible for 10-qubit circuits, although complete schedules cannot be reliably generated.
For the circuit cnt3-5_179, no model produced a valid schedule, and the number of executed gates before failure further decreased compared to mini_alu_305. LLaMA executed 16 gates at $T=1.0$ and 12 gates at $T=0.7$ before failing, while all other models failed after executing between four and nine gates. These results further highlight that shuttling schedule generation becomes increasingly challenging as the number of qubits grows.
Overall, our results show that fine-tuned LLMs can generate valid shuttling schedules and can, in some cases, outperform classical compilers. However, producing complete schedules becomes increasingly challenging as the number of qubits and gates increases. The models frequently enter invalid states, requiring retries that substantially inflate token counts without contributing to the final output. Generalization to previously unseen architectures remains limited. Across all experiments, LLaMA demonstrated the most stable performance, whereas Qwen, DeepSeek, and Gemma failed more frequently and at earlier stages as circuit size increased. When comparing the both Qwen models, it can be noticed that the smaller model with 4 billion parameters produced better results than the the larger model with 14 billion parameters. This indicates, together with LLaMA which is with 3 billion parameters the smallest model we tested, that smaller models can be better adapted to shuttling schedule generation than larger ones.
5 Conclusion and Outlook
We have presented a shuttling compiler for trapped-ion quantum computers based on large language models (LLMs). We fine-tuned several LLMs on datasets derived from shuttling schedules generated by classical compilers for linear and branched one-dimensional trap architectures and evaluated them on circuits of varying sizes. For small circuits on both linear and junction layouts, the models were able to generate valid schedules and in some cases even required fewer operations than the classical compilers.
As circuit size increased, generating complete shuttling schedules became substantially more difficult. Larger numbers of qubits make trap reconfiguration increasingly complex and often require long, recursive sequences of shuttling operations. As a result, the models frequently succeeded in generating valid operations only for an initial portion of the circuit before reaching trap configurations from which continuation would require operations beyond their generation capabilities. Since our compiler constructs schedules gate by gate, the likelihood of entering such unrecoverable configurations grows with circuit depth. These results indicate that LLM-based shuttling compilation is feasible in principle but remains limited in robustness. Evaluations on previously unseen trap layouts further revealed a sharp performance degradation, indicating that the models generalize poorly beyond the architectures encountered during training.
To further improve the capabilities of our compiler, we plan to enhance the training pipeline in two stages: first using Direct Preference Optimization (DPO) [RafailovSMEMF23], and later using Gradient Regularized Policy Optimization (GRPO) [OpenAI24]. Applying DPO requires generating multiple candidate shuttling operations at each step so that the model can learn which alternatives are preferable. As discussed in Sec. 3.3, the schedule with the lowest number of operations cannot always be used as the preferred example, since following it naively may lead the model into local minima that require additional operations to escape. Instead, DPO could help the model identify invalid operations by consistently deprioritizing them. To refine this further, a lookahead strategy would be needed: for each valid candidate, the shuttlings for a certain number of next gates must be computed to determine which initial choice leads to the lowest overall number of operations. While the runtime of this approach grows exponential, it provides valuable multi-step supervision.
Building on this, GRPO offers a framework for reinforcement learning with human feedback (RLHF) [ChristianoLBMSD17]. This would require a continuous reward function that penalizes invalid schedules and rewards valid schedules with fewer operations, again ideally using a lookahead-based evaluation. Due to the large number of shuttling generations needed for DPO and the additional computational overhead of GRPO, scaling these methods will require access to more GPUs than the eight H100 GPUs available for the present work.
Overall, our work establishes the first foundation for LLM-based shuttling compilers for trapped-ion quantum computers and identifies promising directions to significantly extend their performance and reliability in future research.
Acknowledgments
We acknowledge funding by the German Federal Ministry of Education and Research (BMBF) via the VDI within the project IQuAn. We thank the data center (ZDV) of the Johannes Gutenberg University for the use of their Mogon NHR and Mogon KI clusters for evaluating our LLMs.

### v2

@@
 Abstract
-Trapped-ion quantum computers based on segmented traps rely on shuttling operations to establish connectivity between multiple sub-registers within a quantum processing unit.
-Several architectures of increasing complexity have already been realized, including linear arrays, racetrack loops, and junction-based layouts.
-As hardware capabilities advance, the need arises for flexible software layers within the control stack to manage qubit routing—the process of dynamically reconfiguring qubit positions so that all qubits involved in a gate operation are co-located within the same segment.
-Existing approaches typically employ architecture-specific heuristics, which become impractical as system complexity grows.
+Trapped-ion quantum computers based on segmented traps rely on shuttling operations to establish long-range connectivity between sub-registers.
+Qubit routing dynamically reconfigures qubit positions so that all qubits involved in a gate operation are co-located within the same segment, a task whose complexity increases with system size.
 To address this challenge, we propose a layout-independent compilation strategy based on large language models (LLMs).
 Specifically, we fine-tune pretrained LLMs to generate the required shuttling operations.
-We evaluate this approach on both linear and branched one-dimensional architectures, demonstrating that it provides a foundation for developing LLM-based shuttling compilers for trapped-ion quantum computers.
+We evaluate this approach on linear and branched one-dimensional architectures using quantum circuits of up to 16 qubits.
+Our results show that the fine-tuned LLMs generate valid shuttling schedules and, in some cases, outperform previous shuttling compilers by requiring approximately 15 % less shuttle overhead.
+However, results degrade as the algorithms increase in width and depth.
+In future, we plan to improve LLM-based shuttle compilation by enhancing our training pipeline using Direct Preference Optimization (DPO) and Gradient Regularized Policy Optimization (GRPO).
 1 Introduction
-Shuttling-based trapped-ion quantum computers [KielpinskiMW02, KausahlLSHPSBMSP20, PinoDFGMABFHMRN21, MuraliDBM20, MosesB+23] are currently limited by both qubit count and the depth of reliably executable quantum circuits.
-Recent advances are gradually overcoming these barriers, enabling demonstrations of quantum advantage [AruteAB+19] and opening pathways toward practical applications in real-world computational tasks [PeruzzoMSYZLAO14, AnschuetzOAC19].
-At the same time, this progress introduces additional hardware complexity, necessitating more sophisticated compilation strategies for efficient qubit routing and quantum circuit execution.
-In trapped-ion systems, atomic ions are confined in segmented microchip traps [SchulzPSS06].
-Each pair of electrodes generates a localized potential well that can hold multiple ions, effectively partitioning the trap into smaller, more manageable segments.
-The quantum information is encoded in the internal states of individual ions, each of which represents a qubit; accordingly, we use the term qubit to refer to an ion throughout the paper.
-By adjusting electrode voltages, trapping potentials can be dynamically reconfigured, enabling qubits to be shuttled between segments for successive gate operations.
-Segments are typically categorized as either storage segments, where qubits are held, or gate segments, where quantum operations are performed using laser or microwave radiation.
-In accordance with the current maturity level of trapped-ion hardware, the number of available gate segments is still very limited and often restricted to a single one [KausahlLSHPSBMSP20, SchmaleTBPKDOWB22].
-Most current architectures arrange trap segments in a linear array [KausahlLSHPSBMSP20, PinoDFGMABFHMRN21, BowlerGLTHJHLW12, LeeJPJKC21] or in a racetrack loop [MosesB+23], where each segment connects to two neighbors.
+Shuttling-based trapped-ion quantum computers [28, 27, 47, 41, 40] are currently limited by the number of qubits and the quality of gate operations.
+We are addressing the complexity that arises when a large number of qubits is routed efficiently for the execution of quantum algorithms.
+For a scalable shuttle-based approach of a quantum computer, the quantum information is encoded in the internal states of individual ions, each of which represents a qubit; accordingly, we use the term qubit to refer to an ion throughout the paper.
+The qubits are confined in segmented microchip traps [62], where each segment generates a localized potential well.
+By adjusting electrode voltages, trapping potentials can be dynamically reconfigured, enabling qubits to be shuttled between segments.
+Segments are either storage segments, where qubits are held, or gate segments.
+When two qubits are in the same gate segment, laser- or microwave-driven gate operations are executed.
+In accordance with the current maturity level of trapped-ion hardware, the number of available gate segments is still very limited [27, 57].
+In this work we restrict ourselves to a single gate segment, while the approach can be extended to a number of gate segments, which have been realized recently [40, 52].
+Current architectures arrange trap segments in a linear array [27, 47, 5, 33] or in a racetrack loop [40], where each segment connects to two neighbors.
 Such designs limit scalability, as the number of qubits scales linearly with the physical size of the trap.
-Introducing junctions [HensingerOSHYADMR06, BlakestadOVABLW09, WrightAFVDHPLDKSH13, ShuVBNVSB14] increases connectivity, allowing segments to connect to multiple neighbors.
-Quantinuum has demonstrated this idea in their Helios architecture, where a ring-shaped storage region is connected to two linear legs via an $\mathsf{X}$-shaped junction [Ransford+25].
+Introducing junctions [24, 4, 75, 64] increases connectivity, allowing segments to connect to multiple neighbors.
+Quantinuum has demonstrated this idea in their Helios architecture, where a ring-shaped storage region is connected to two linear legs via an $\mathsf{X}$-shaped junction [52].
 However, Helios contains only a single junction.
-Increasing the number of junctions would enable more complex topologies in the future, such as two-dimensional grids [LekitschWFMDWH17].
-Executing a quantum circuit on such hardware requires a sequence of physical operations.
-Because quantum gates can only be applied to qubits within the same gate segment, the trap must be dynamically reconfigured using four fundamental shuttling operations:
-  • Translation: Moves the qubits from one trap segment to a neighboring segment.
-  • Separation: Splits the qubits stored in one segment into two segments.
-  • Merge: Combines qubits from two segments into a single segment.
-  • Swap: Physically rotates the content of a segment by 180 degrees, reversing the order of the qubits within the segment.
-While shuttling operations enable effective all-to-all connectivity between qubits and strongly suppress cross-talk errors, they also introduce substantial time overhead that lengthens operation timescales [HilderPOSOLRMSP22].
-As a result, error sources such as dephasing and noise-induced heating become more pronounced, while the shuttling process itself can further degrade gate fidelities through ion heating.
-The gate fidelity quantifies how closely the implemented quantum gate matches its ideal operation, with lower fidelities indicating higher error probabilities during gate execution.
+Increasing the number of junctions would enable more complex topologies in the future, such as a two-dimensional quantum CCD [34].
+Executing a quantum circuit requires a sequence of physical operations. The trap must be dynamically reconfigured using four fundamental shuttling operations:
+  • Translation: Moves the qubits to a neighboring segment.
+  • Separation: Splits two qubits from one segment into two different segments.
+  • Merge: Joins one qubit from the left and one from the right into one segment.
+  • Swap: Physically rotates within one segment, thus reversing the order of the qubits.
+Shuttling operations introduce substantial time overhead [25].
+As a result, error sources such as dephasing and noise-induced heating become more pronounced, and the shuttling process itself can further degrade gate fidelities through ion motional excitation and heating.
 Consequently, the number of shuttling operations must be minimized, highlighting the need for efficient compilation routines.
-Compilation refers to the translation of a logical quantum circuit, which comprise abstract qubits and idealized gates, into a sequence of hardware-level instructions executable on a specific trapped-ion device.
-This process includes decomposing logical gates into the device’s native gate set, mapping logical to physical qubits, and determining a shuttling schedule that defines an executable sequence of shuttling operations.
-The resulting schedule must adhere to hardware constraints while minimizing resource overheads such as operation time, qubit transport distance, and cumulative error.
-Existing shuttling compilers compute valid qubit routing schedules by generating shuttling schedules tailored to specific architectures, such as linear arrays [PinoDFGMABFHMRN21, Wagner22, DurandauWMBSPB23], racetrack loops [MosesB+23], drive-through architectures [ChangJCHL25], plus-sign layouts [SchmaleTBPKDOWB22], a ring with linear legs [Ransford+25], branched one-dimensional traps [KreppelMWHPSB24], and two-dimensional designs [WebberHWH20, DaiBR24, SchoenbergerHBW25, SchoenbergerHSW25, SchoenbergerW25].
-However, as architectures grow more complex, designing separate heuristics for each layout becomes increasingly impractical.
-To overcome the limitations of architecture-specific compilation, large language models (LLMs) offer a promising, layout-independent alternative.
-LLMs can perform new tasks from textual instructions or a few demonstrations [Brown+20], making them well suited for learning shuttling behavior directly from examples.
-In this work, we investigate whether current LLMs can generate valid and efficient shuttling schedules and identify the challenges that arise when applying them to trapped-ion shuttling.
-We fine-tune LLaMA 3.2 [MetaAI24], Qwen 3 [QwenTeam25], DeepSeek LLM [Deepseek-AI24], and Gemma 3 [GemmaTeam25] on datasets derived from classical shuttling compilers for both linear [Wagner22, DurandauWMBSPB23] and branched one-dimensional architectures with junctions [KreppelMWHPSB24].
-We then compare their performance across multiple benchmark circuits and additionally evaluate their ability to generalize to previously unseen trap layouts.
-Our evaluation shows that the fine-tuned models were able to generate valid shuttling schedules and, in some cases, even surpassed classical shuttling compilers.
-However, performance degraded as the number of qubits and gates increased: all models frequently reached trap configurations that require numerous retries or prevent further schedule generation.
-Results on the unseen trap layouts further indicate that the models’ ability to generalize beyond the training architectures remains limited.
-Overall, the results show that our approach provides a solid foundation for developing LLM-based shuttling compilers.
-Our compiler differs from machine-learning-based routing techniques developed for superconducting architectures [PalerSFA23, PozziHSM22, FanGL22, PascoalFA24, TangDKFKS24, SinhaAS22, RussoPPAC25, SangHH25, ZeynaliB25, CuomoCKTAPC23, EscofetOBPVFAAA25, SundaramGR25].
+Compilation refers to the translation of a quantum circuit, which comprises abstract qubits and idealized gates, into a sequence of instructions executable on a specific trapped-ion device.
+This process starts out by decomposing gates into the device’s native gate set, then mapping qubits to trapped ions, and finally determining a shuttling schedule that defines an executable sequence of shuttling operations.
+The resulting schedule must adhere to the architecture of the processor while minimizing resource overheads such as operation time, qubit transport distance, etc.
+Existing shuttling compilers compute valid qubit routing schedules by generating shuttling schedules tailored to specific architectures [47, 71, 19, 40, 7, 57, 52, 30, 72, 14, 59, 58, 61].
+Due to the architectures growing complexity, designing separate heuristics for each layout becomes increasingly impractical.
+We investigate large language models (LLMs) as a layout-independent alternative.
+LLMs can perform new tasks from textual instructions or a few demonstrations [6], making them well suited for learning shuttling behavior directly from examples.
+In this work, we examine whether current LLMs can generate valid and efficient shuttling schedules and identify the challenges that arise when applying them to trapped-ion shuttling.
+We fine-tune LLaMA 3.2 [37], Qwen 3 [49], DeepSeek LLM [16], and Gemma 3 [22] on datasets derived from classical shuttling compilers for both linear [71, 19] and branched one-dimensional architectures with junctions [30].
+We then compare the performance across multiple benchmark circuits.
+Moreover, we evaluate the ability to generalize to novel trap layouts.
+Our compiler differs from machine-learning-based routing techniques developed for superconducting architectures [45, 48, 21, 46, 68, 65, 55, 56, 77, 13, 20, 67].
 While those methods primarily optimize SWAP placement and gate parallelism on fixed coupling graphs, our approach accounts for the hardware-specific characteristics of trapped-ion devices.
-In particular, it considers shuttling operations, which are absent in superconducting systems, and the limited number of gate segments, which restrict the parallel execution of quantum gates.
-Our work is the first to employ large language models for compiling quantum circuits into shuttling schedules for trapped-ion quantum computers and to compare different LLMs for this task.
+Our work is the first to employ LLMs for compiling quantum circuits into shuttling schedules for trapped-ion quantum computers and to compare different LLMs for this task.
 The remainder of this paper is structured as follows: Sec.
@@
 […]
-Within these compilation stacks, the individual compilation stages can be realized using a variety of algorithmic techniques.
-Compilation typically begins with a transformation to the hardware’s native gate set.
-Different methods pursue different optimization goals, such as reducing circuit depth, minimizing the number of specific gates, or adapting to particular architectures.
-Some compilers apply heuristics to obtain approximate solutions [YounisI22, JinHCHZZ23, KreppelMOWHPSB23, AroraXWLLDA25], while others employ SAT solvers to produce optimal results for smaller instances [PehamBKWB23].
-More recently, machine learning has been applied to this task, ranging from recommender models [WeidenYKKI23] to reinforcement learning agents trained for specific optimization goals [MoroPRP21, FöselNML+22, Preti24, RietschDUPPMS24, Ruiz+25, RiuNVGE25, WeidenKKI25].
-After this transformation, logical qubits must be mapped to physical qubits, and routing must be computed to enable the execution of gates under architectural constraints.
-This step is architecture-dependent, and techniques have been developed for a variety of platforms.
-For example, methods exist for neutral atom architectures [SchmidPW24, NottinghamPSWBCB24, StadeSVW24] and spin-qubit systems [ParaskevopoulosSAF23, ParaskevopoulosAF24], but most work focuses on superconducting and trapped-ion platforms.
-In superconducting devices, architectures are modeled as coupling graphs, where vertices represent qubits and edges denote possible two-qubit interactions.
-Since two-qubit gates are restricted to adjacent qubits, additional SWAP gates are required to establish long-range connectivity.
-At the same time, gates on independent qubits can be executed simultaneously.
-Existing approaches differ in their optimization goals, such as minimizing SWAPs or maximizing gate parallelism.
-Proposed techniques span a wide range of methods, including heuristics and mathematical formulations [MaslovFM08, SaeediWD11, ShafaeiSP14, WilleLD14, LinSJ15, PedramS16, SiraichiSCP18, HattoriY18, LiDX19, MuraliBJCM19, ChildsSU19, CowtanDDKSS19, MatsuoY19, ZulehnerBW19, ZulehnerPW19, ItokoRIM20, NiuSST20, AlamAG20, TanC20, ZhouLF20, ZhangHQJYZ21, LiWSJDX22, ParkKKSK22, NanniciniBGJ22, HuaJCVKBLJZ23, PehamBW23, KhandavilliPNLND23, LiuYWHKI23, FuZWXSL23, ShaikP23, YangKSHD24, ChengYKWCH24, KattemölleH25], as well as machine-learning-based approaches.
-The latter include k-nearest neighbors, random forests, and multilayer perceptrons for qubit mapping [PalerSFA23], and reinforcement learning (RL) methods for routing [PozziHSM22, FanGL22, PascoalFA24, TangDKFKS24].
-Specifically, [TangDKFKS24] employs a transformer-based neural network as an RL agent and reports a reduction in the number of SWAP gates by 10 to 20 % compared to the SABRE heuristic [LiDX19].
-Hybrid approaches combining reinforcement learning with graph neural networks (GNNs) have also been proposed [SinhaAS22].
-Beyond single-device architectures, several works address distributed superconducting architectures, where qubits are partitioned across multiple devices and minimizing inter-device qubit transfers becomes critical [RussoPPAC25, SangHH25, ZeynaliB25, CuomoCKTAPC23, EscofetOBPVFAAA25, SundaramGR25].
-In this setting, GNN- and transformer-based models are used to capture global circuit structure [RussoPPAC25, SangHH25], and qubit reuse strategies are explored to further reduce communication overhead [SangHH25].
-These approaches demonstrate reductions of up to 100 % in inter-device qubit transfers, effectively eliminating them when qubit reuse is enabled, and reductions of up to 99 % compared to QUBO [BandicPNORAVVAAF23] even without qubit reuse.
-The improvements are attributed to RL-based methods learning long-range dependencies across the entire circuit, enabling globally optimized qubit placements.
-Compared to Qiskit [WilleMN19], however, the reduction in inter-device transfers is in the range of 15 to 40 %, and some circuits still incur higher communication costs.
-Finally, [ZeynaliB25] proposes a noise-adaptive mapping strategy for distributed architectures that combines reinforcement learning for qubit placement with long short-term memory (LSTM) networks for noise migration and a multi-head self-attention mechanism to identify long-range qubit dependencies.
-Experimental results show that this approach increases the mean circuit fidelity by 49 %, raising it from 0.62 to 0.92, while reducing inter-chip operations by an average of 80 % compared to QUBO.
-Trapped-ion quantum computers exhibit distinct hardware characteristics.
-In these systems, multiple qubits are confined together within individual trap segments.
-While quantum gates can only be applied at specific segments, the qubit distribution across segments can be dynamically reconfigured through shuttling operations.
-Shuttling compilers compute sequences of such operations, enabling circuit execution on a given trap.
-Starting from a quantum circuit, they map logical qubits to physical ones and then derive shuttling schedules.
-While the minimal number of timesteps can be computed using Boolean satisfiability methods [SchoenbergerHBW24], most existing compilers employ heuristics to reduce the number of shuttling operations.
-Prior work includes design studies analyzing how architectural choices affect shuttling overhead [MuraliDBM20, AshSakiTG22], as well as compilers targeting specific trap configurations as mentioned in Sec.
+Due to the hardware characteristics of shuttling-based trapped-ion systems, shuttling compilers are a crucial component of the quantum compilation stack.
+Starting from a quantum circuit, they map qubits to ions and then derive shuttling schedules.
+While the minimal number of timesteps can be computed using Boolean satisfiability methods [60], most existing compilers employ heuristics to reduce the number of shuttling operations.
+Prior work includes design studies analyzing how architectural choices affect shuttling overhead [41, 1], as well as compilers targeting specific trap configurations [47, 71, 19, 40, 7, 57, 52, 30, 72, 14, 59, 58, 61].
+However, all these existing approaches are tightly coupled to specific architectures.
+Beyond single-trap systems, compilation has also been explored for distributed trapped-ion architectures, where multiple traps are connected via photonic interconnects that mediate quantum communication between nodes [76].
+Such systems introduce additional scheduling and routing challenges, further highlighting the need for flexible compilation strategies that generalize across architectures.
+3 Compiler
+This section explains how we fine-tune LLMs to generate shuttling schedules for quantum circuit execution. The process comprises three phases: dataset generation, fine-tuning, and inference.
+3.1 Dataset Generation
+First, we require examples that illustrate how shuttling is performed.
+These examples are provided as shuttling schedules, as shown in Fig.
 1.
-However, all these existing approaches are tightly coupled to specific architectures.
-Beyond single-trap systems, compilation has also been explored for distributed trapped-ion architectures, where multiple traps are connected via photonic interconnects that mediate quantum communication between nodes [WuZWW25].
-Such systems introduce additional scheduling and routing challenges, further highlighting the need for flexible compilation strategies that generalize across architectures.
-In contrast to prior work, we present the first layout-independent shuttling compiler for trapped-ion quantum computers based on large language models (LLMs) and provide a comparison of different models for this task.
-By fine-tuning on examples of shuttling problems, LLMs are able to generalize beyond a single trap design and generate valid and efficient shuttling schedules for a variety of architectures, provided appropriate training data is available.
-3 Compiler
-This section explains how we fine-tune LLMs to generate shuttling schedules for quantum circuit execution.
-The process comprises three phases, illustrated in Fig.
-1 and detailed in the following subsections.
-Fig.
-1: Overview of the three phases of our compiler.
-We first generate training and evaluation datasets in Alpaca format, where each entry consists of an instruction and its corresponding output.
-These datasets are then used to fine-tune an LLM with Axolotl, which involves pre-processing and training.
-Finally, the fine-tuned model is deployed for inference to compute shuttling schedules by submitting instructions to a vLLM server.
-3.1 Dataset Generation
-To fine-tune an LLM, we require examples that illustrate how shuttling is performed.
-These examples are provided as shuttling schedules, as shown in Fig.
-2.
 Since the physical movements depend only on which qubits participate in each gate, rather than on the specific gate function, the gate functionality is omitted in this representation.
-Such schedules can be generated in different ways, for example by constructing random sequences, by execution guided through a reward function, or by employing another (not necessarily optimized) compiler.
+Schedules can be generated in different ways, for example by constructing random sequences, by execution guided through a reward function, or by employing another (not necessarily optimized) compiler.
 The number of required schedules depends on the trap architecture, its size, and the number of qubits.
@@
 Fig.
-2: Shuttling schedule of a quantum circuit with two qubits and three gates on a linear architecture consisting of three segments.
+1: Shuttling schedule of a quantum circuit with two qubits and three gates on a linear architecture consisting of three segments.
 Both qubits are initially located in gate segment 11.
@@
 In this work, we focus on linear architectures and branched one-dimensional architectures with junctions, for which example graphs are shown in Fig.
-3.
+2.
 (a) Graph representation of a linear architecture with 15 segments.
@@
 Fig.
-3: Graph representations of ion-trap architectures.
+2: Graph representations of ion-trap architectures.
 Red vertices denote gate segments (GS), yellow vertices denote storage segments, and green vertices denote junctions, which are the only vertices connected to more than two others.
@@
 Both datasets consist of multiple data entries, obtained by disassembling the shuttling schedules as illustrated in Fig.
-2.
+1.
 Each entry corresponds to the portion of the schedule between two gate executions.
@@
 When constructing datasets, it is also possible to mix schedules from different architectures so that a single LLM can be fine-tuned across them.
-Each data entry follows the Alpaca format [TaoriGZDLGLH23], consisting of an instruction and an output. Since the LLM can only process textual input, all required information is encoded as text.
+Each data entry follows the Alpaca format [69], consisting of an instruction and an output. Since the LLM can only process textual input, all required information is encoded as text.
 3.1.1 Instruction
@@
 Each qubit is assigned a position within this chain, allowing it to be uniquely referenced as a tuple $[v,p]$, where $v$ denotes the vertex number and $p$ the qubit’s position in the chain.
-Fig.
-4: Quantum circuit with four qubits and 14 gates.
-Purple gates form the first layer and are executable when their qubits are located in a gate segment, whereas yellow gates belong to deeper layers and are not yet executable.
-For each qubit, gates are numbered in ascending order to indicate their execution sequence.
 In addition, the instruction specifies the quantum gates of the circuit, which must be executed on the given architecture.
@@
 Due to architectural constraints, a gate is executable only if all involved qubits reside in the same gate segment, and this segment must not contain any additional qubits.
-As illustrated in Fig.
-4, the gates are divided into two groups: gates in the first layer of the circuit which are executable when the corresponding qubits are in a gate segment and not yet executable gates in deeper layers.
+The quantum circuit can be partitioned into layers such that the first layer contains only gates whose involved qubits have no predecessors and are executable once the corresponding qubits are in the gate segment, while gates in deeper layers have at least one qubit with a predecessor and are therefore not yet executable.
+This layer-based division is illustrated in Fig.
+3.
 For each qubit, the gate numbers are assigned in ascending order.
@@
 Finally, to link the circuit qubits with the qubits in the trap, all qubits in the gate descriptions are replaced by tuples $[v,p]$ of their corresponding position in the trap.
+Fig.
+3: Quantum circuit with four qubits and 14 gates.
+Purple gates form the first layer and are executable when their qubits are located in the gate segment, whereas yellow gates belong to deeper layers and are not yet executable.
+For each qubit, gates are numbered in ascending order to indicate their execution sequence.
 To modify the trap state, the LLM requires a description of the available shuttling operations.
@@
 In segmented trapped-ion architectures, junctions are formed by the intersections of electrodes, enabling qubits to be routed between different linear sections.
-Unlike storage segments, junctions do not provide a stable trapping potential suitable for holding qubits.
-Instead, the electrode potentials are shaped to guide qubits smoothly through the junction, ensuring efficient transport.
-Consequently, qubits must pass through junctions without stopping, and reversing direction within a junction is infeasible because the electric fields are engineered for unidirectional traversal rather than oscillatory motion [WrightAFVDHPLDKSH13, MokhberiSW17].
-When computing shuttling operations, the LLM must respect this physical constraint.
+Qubits must pass through junctions without stopping, and reversing direction within a junction is infeasible because the electric fields are engineered for unidirectional traversal rather than oscillatory motion [75, 39].
 In particular, Separate, Merge, Swap, and gate execution cannot be performed on junction vertices, which must remain empty throughout execution.
@@
   • Swap: Let $v\in\mathcal{V}$ be a vertex eligible for swapping, containing at least two qubits. Reverse the order of the qubit chain in $v$.
-  • Execute Gate: Perform one of the first-layer gates in a gate segment, provided all previously defined execution conditions are satisfied.
+  • Execute Gate: Perform the quantum gate in the gate segment, once all previously defined execution conditions are satisfied.
 After specifying all rules to the LLM, the overall goal is defined: execute all gates using the minimal number of operations.
@@
   2. 2)
-    First-layer gates: All gates in the first layer of the circuit that are executable when their corresponding qubits reside in a gate segment, along with the tuples of the qubits they act on.
+    First-layer gates: All gates in the first layer of the circuit that are executable when their corresponding qubits reside in the gate segment, along with the tuples of the qubits they act on.
   3. 3)
@@
 The instruction concludes by defining the task for the LLM: generate a shuttling sequence that enables the execution of one of the first-layer gates, contributing to the overall goal.
-After each operation, the LLM should also provide the updated qubit positions, the list of first-layer gates with updated position tuples of the qubits, and the new set of allowed shuttling operations.
+After each operation, the LLM should also provide the updated qubit positions, the list of first-layer gates with the updated position tuples of the qubits, and the new set of allowed shuttling operations.
 Requiring these additional outputs at each step supports the LLM to reason carefully about the current state and the implications of each operation.
 3.1.2 Output
-The task of the LLM is to generate a shuttling sequence which allows to execute one of the first-layer gates.
-To learn the generation of such sequences, the LLM needs examples of expected shuttling sequences during the training phase that solve the task given in the instruction.
+To learn, the LLM needs examples of expected shuttling sequences during the training phase that solve the task given in the instruction.
 Therefore the output contains the part of the previous generated shuttling schedule that shows how one of the first-layer gates can be executed from the given state.
-As long as not all requirements are fulfilled to execute one of the executable gates, usually multiple shuttling operations are required in order to execute one of these gates.
-The number of operations depend on the size of the trap and the number of qubits used.
+As long as not all requirements are fulfilled to execute one of the gates, usually multiple shuttling operations are required.
+The number of operations depends on the size of the trap and the number of qubits used.
 The reason is that with increasing trap size and qubit number, more shuttling operations are required to reconfigure the trap segments so that the trap is in a valid state for a gate execution.
-The output begins with the first shuttling operation which is to be executed based on the state given in the instruction.
-Since the instruction demands also to output the updated qubit positions and based on this the first-layer gates and the new allowed shuttling operations after each operation, these information is also contained and presented in the same way as in the instruction.
-This information helps the LLM to understand which changes happen in the trap state based on the executed shuttling operation and which consequences they have.
+The output begins with the first shuttling operation which is to be executed.
+This includes the updated qubit positions, the first-layer gates, and the new allowed shuttling operations.
+This information helps the LLM to learn which changes happen with the executed shuttling operation.
 Afterwards, the output contains the next shuttling operation based on the new state.
-This continues until the operation “Execute Gate” indicates that one of the first-layer gates is to be executed.
-After generating the training and evaluation datasets, we use them to fine-tune an LLM with Axolotl [Axolotl23].
+This continues until the operation “Execute Gate” indicates that the ions are arranged such that a first-layer gate can be executed.
 3.2 Fine-tuning with Axolotl
-After dataset preparation, we fine-tune pre-trained LLMs from Huggingface [Wolf+20] to compute shuttling schedules.
-For this, we employ Axolotl [Axolotl23], which provides seamless compatibility with Huggingface models.
-Axolotl supports various training approaches, including full fine-tuning, Low-Rank Adaptation (LoRA) [HuSWALWWC22], and Quantized LoRA (QLoRA) [DettmersPHZ23].
-Since LoRA and QLoRA did not yield the expected results (see Sec.
+After dataset preparation, we fine-tune pre-trained LLMs from Huggingface [74] to compute shuttling schedules.
+For this, we fine-tune with Axolotl [2], which provides seamless compatibility with Huggingface models.
+Axolotl supports various training approaches, including full fine-tuning, Low-Rank Adaptation (LoRA) [26], and Quantized LoRA (QLoRA) [18].
+Since LoRA and QLoRA did not yield good results (see Sec.
 3.4), we performed full fine-tuning of the LLM.
-Furthermore, Axolotl offers performance optimizations such as multipacking [KunduLWGM24], FlashAttention 2 [Dao24], sequence (context) parallelism [NarayananSCLPKVKBCPZ21], and multi-GPU training with DeepSpeed [RasleyRRH20], which we describe in more detail below.
+Furthermore, Axolotl offers performance optimizations such as multipacking [31], FlashAttention 2 [15], sequence (context) parallelism [42], and multi-GPU training with DeepSpeed [53], which we describe in more detail below.
 3.2.1 Performance Optimizations
 To minimize resource requirements and enable fine-tuning on as few GPUs as possible, it is essential to reduce memory usage and parallelize execution.
-We achieve this with DeepSpeed, employing the Zero Redundancy Optimizer (ZeRO) at optimization level 3 [RajbhandariRRH20] to accelerate and distribute fine-tuning [ShoeybiPPLCC19].
+We achieve this with DeepSpeed, employing the Zero Redundancy Optimizer (ZeRO) at optimization level 3 [51] to accelerate and distribute fine-tuning [63].
 In this setting, 16-bit model parameters are partitioned across GPUs, while DeepSpeed manages parameter partitioning and collection during forward and backward passes.
 To further reduce memory consumption, we use bfloat16 as the 16-bit floating-point format.
-Additionally, we apply ZeRO Offloading [RenRARYZLH21], which transfers optimizer states, as well as model parameters and gradients when inactive, to the host CPU and RAM.
+Additionally, we apply ZeRO Offloading [54], which transfers optimizer states, as well as model parameters and gradients when inactive, to the host CPU and RAM.
 The data entries in our datasets vary considerably in length due to differing output sizes, which depend on the number of shuttling operations required to execute one of the first-layer gates.
@@
 For cases where individual sequences exceed the VRAM capacity of a single GPU, we use sequence parallelization.
-This approach relies on blockwise ring attention [LiuZA24], which enables distributed computation of self-attention and feedforward layers across multiple GPUs, while overlapping communication and computation to minimize overhead.
+This approach relies on blockwise ring attention [35], which enables distributed computation of self-attention and feedforward layers across multiple GPUs, while overlapping communication and computation to minimize overhead.
 Moreover, because each instruction in our dataset specifies a problem and its context, while the corresponding output provides a step-by-step solution, we compute the loss exclusively on the output.
@@
 We use gradient accumulation over eight batches to mitigate gradient noise and to improve convergence despite the small batch size.
-The accumulated gradients are applied by the AdamW optimizer [LoshchilovH19] to update the network weights.
+The accumulated gradients are applied by the AdamW optimizer [36] to update the network weights.
 In addition to training, the Trainer handles logging, evaluation at regular intervals, and checkpointing.
@@
 3.3 Inference with vLLM
-After fine-tuning, the model is deployed for inference using vLLM [KwonLZSZYGZS23], which provides efficient multi-GPU support.
+After fine-tuning, the model is deployed for inference using vLLM [32], which provides efficient multi-GPU support.
 Its PagedAttention mechanism optimizes memory management of attention keys and values, thereby reducing overhead.
@@
 While such a lookahead strategy could reduce the overall number of operations, its runtime would grow exponentially with the number of steps.
+Fig.
+4: Overview of the three phases of our compiler.
+We first generate training and evaluation datasets in Alpaca format, where each entry consists of an instruction and its corresponding output.
+These datasets are then used to fine-tune an LLM with Axolotl, which involves pre-processing and training.
+Finally, the fine-tuned model is deployed for inference to compute shuttling schedules by submitting instructions to a vLLM server.
+In summary, the complete compilation flow of our compiler is illustrated in Fig. 4.111The datasets, training scripts, and trained models are available from the authors upon request.
 3.4 Alternative Approaches and Challenges
@@
 For the example in Fig.
-2, this would result in nine entries instead of three.
+1, this would result in nine entries instead of three.
 While this approach reduced the size of individual entries, it substantially increased their total number.
@@
 Although the chosen operations were valid, they did not necessarily advance the execution of the quantum circuit.
-Second, instead of limiting the instruction to the first-layer gates and a subset of deeper-layer gates, we included the entire remaining circuit in OpenQASM format [CrossBSG17, CrossJADBHRSSGJ22], along with a mapping from qubits to their positions in the trap.
+Second, instead of limiting the instruction to the first-layer gates and a subset of deeper-layer gates, we included the entire remaining circuit in OpenQASM format [12, 11], along with a mapping from qubits to their positions in the trap.
 While this approach has much longer state descriptions, it also introduced a critical drawback: the LLM was unable to reliably identify the first-layer gates and frequently generated shuttling operations that targeted arbitrary, and often deeper-layer, gates.
@@
 5: Graph representations of ion-trap architectures used to evaluate the generalization capabilities of the fine-tuned LLMs on previously unseen layouts.
-Architectures (a) and (b) feature three-way junctions, while architecture (c) contains four-way junctions.
-For each layout, the number of yellow storage vertices was scaled to match the qubit count of the respective test circuit.
+Architectures (a) and (b) feature three-way junctions, while architecture (c) contains four-way junctions (green).
+For each layout, the number of storage vertices (yellow) was scaled to match the qubit count of the respective test circuit.
+GS (red) indicates the segment, where quantum gates are executed.
 4.1 Setup
-In the following, we first describe the experimental setup used for fine-tuning the models and then outline the setup employed for inference with the fine-tuned LLMs.
+In the following, we first describe the setup used for fine-tuning the models and then outline the setup employed for inference with the fine-tuned LLMs.
 4.1.1 Fine-tuning
@@
 For the linear architecture, the number of qubits was varied between 2 and 16.
-For the junction architecture, the number of qubits ranged from 3 to 16, and we additionally varied the stack depth, specifying the number of storage segments per stack, and the junction distance, defining the spacing between the junctions on each side of the gate segment, as illustrated in 3(b).
+For the junction architecture, the number of qubits ranged from 3 to 16, and we additionally varied the stack depth, specifying the number of storage segments per stack, and the junction distance, defining the spacing between the junctions on each side of the gate segment, as illustrated in 2(b).
 In the following, we denote the stack depth by $s$ and the junction distance by $d$.
 Due to hardware limitations, we restricted the datasets to circuits of these sizes, as larger circuits would require more complex shuttling patterns and, consequently, additional VRAM and GPUs for fine-tuning.
-Shuttling schedules were generated using the compilers from [Wagner22] for the linear architecture and [KreppelMWHPSB24] for the branched architecture.
-The underlying quantum circuits were randomly generated with Qiskit [WilleMN19], containing single- and two-qubit gates with a depth of five.
-These circuits were then compiled into a trapped-ion-compatible format using the compiler from [KreppelMOWHPSB23].
+Shuttling schedules were generated using the compilers [71] and [30] for the linear and branched architecture, respectively.
+The underlying quantum circuits were randomly generated with Qiskit [73], containing single- and two-qubit gates with a depth of five.
+These circuits were then compiled into a trapped-ion-compatible format [29].
 For the linear architecture, 120 shuttling schedules per qubit count were included in the training dataset and 30 in the evaluation dataset.
@@
 Both datasets contained approximately equal proportions of examples from each architecture and were automatically shuffled by Axolotl prior to fine-tuning to ensure uniform data distribution.
-Due to the hardware constraints and the size of the dataset entries, model sizes were limited to 50 GB, preventing us from fine-tuning large models such as DeepSeek V3 [Deepseek-AI24_2], DeepSeek R1 [GuoYZ+25], or LLaMA 4 [MetaAI25].
-Within these limits, we fine-tuned a variety of models, including LLaMA 3.2 with 3 billion parameters [MetaAI24], Qwen 3 with 4 and 14 billion parameters [QwenTeam25], DeepSeek LLM with 7 billion parameters [Deepseek-AI24], and Gemma 3 with 27 billion parameters [GemmaTeam25].
-In addition, we attempted to fine-tune GPT-OSS with 20 billion parameters [OpenAI25], which required replacing DeepSpeed with FSDP2 [Zhao+23].
+Due to the hardware constraints and the size of the dataset entries, model sizes were limited to 50 GB, preventing us from fine-tuning large models such as DeepSeek V3 [17], DeepSeek R1 [23], or LLaMA 4 [38].
+Within these limits, we fine-tuned a variety of models, including LLaMA 3.2 with 3 billion parameters [37], Qwen 3 with 4 and 14 billion parameters [49], DeepSeek LLM with 7 billion parameters [16], and Gemma 3 with 27 billion parameters [22].
+In addition, we attempted to fine-tune GPT-OSS with 20 billion parameters [44], which required replacing DeepSpeed with FSDP2 [78].
 However, we were unable to evaluate this model on the test set because it consistently produced invalid responses after fine-tuning: the first one or two shuttling operations were always truncated, making it impossible to construct a valid schedule.
@@
 Due to its larger model size, Gemma was evaluated on a single A100 GPU with 80 GB VRAM on the Mogon KI cluster.
-As test circuits for compilation with the finetuned LLMs, we selected three of the 153 circuits from the library of [Xiangzhen19], which have been used in prior work [ZhouLF20, CowtanDDKSS19, KreppelMOWHPSB23, KreppelMWHPSB24]: 4mod5-bdd_287 (7 qubits, 106 gates), mini_alu_305 (10 qubits, 263 gates), and cnt3-5_179 (16 qubits, 230 gates).
-The gate counts refer to the circuits after compilation using [KreppelMOWHPSB23].
+As test circuits for compilation with the fine-tuned LLMs, we selected three of the 153 circuits from the library of [80], which have been used in prior work [79, 10, 29, 30]: 4mod5-bdd_287 (7 qubits, 106 gates), mini_alu_305 (10 qubits, 263 gates), and cnt3-5_179 (16 qubits, 230 gates).
+The gate counts refer to the circuits after compilation using [29].
 We chose these circuits because they span different qubit counts and, within the library, have the lowest gate counts for their respective qubit numbers.
@@
 As in the training data, each graph contained a single gate segment, and the separate, merge, and swap operations were executable only within that segment.
-We also evaluated the stochastic behavior of the models by varying the temperature $T$, which controls how deterministic the output is.
-Lower temperatures yield more deterministic outputs, with $T=0.0$ making the model behave fully deterministically.
-In contrast, higher temperatures increase the level of randomness in the generated outputs.
+We also evaluated the stochastic behavior of the models by varying $T$, which controls how deterministic the output is.
+Lower $T$ yields more deterministic outputs, and with $T=0.0$ the model becomes fully deterministic.
+Higher $T$ increase the level of randomness in the generated outputs.
 To balance determinism and variability, we evaluated each model–circuit–architecture combination at $T=0.7$ and $T=1.0$.
@@
 Tab.
-1: Operation counts of shuttling schedules generated by LLaMA, Qwen, DeepSeek, and Gemma with different temperatures $T$ compared to the classical compiler.
+1: Operation effort of shuttling schedules generated by LLaMA, Qwen, DeepSeek, and Gemma with different $T$ values compared to the baseline [71, 30].
 For each model, only the run with the fewest operations is reported.
@@
 If all junction configurations failed, the table reports the configuration that completed the most gates before failure.
-Archi- tecture Classical LLaMA 3.2-3B Qwen 3-4B Qwen 3-14B DeepSeek-7B Gemma 3-27B
+Archi- tecture Baseline LLaMA 3.2-3B Qwen 3-4B Qwen 3-14B DeepSeek-7B Gemma 3-27B
 $T=0.7$ $T=1.0$ $T=0.7$ $T=1.0$ $T=0.7$ $T=1.0$ $T=0.7$ $T=1.0$ $T=0.7$ $T=1.0$
@@
 The minimal number of shuttling operations required by the schedules generated by LLaMA, Qwen-4B, Qwen-14B, DeepSeek, and Gemma is summarized in Tab.
-1, together with the corresponding values from the classical compiler.
+1, together with the corresponding values from the baseline compilers [71, 30].
 If a model failed to produce a valid schedule, the table reports how many gates were executed before the generation process became invalid.
@@
 2, including the number of required retries and the number of response tokens.
-We measure the number of generated tokens rather than runtime because token counts are hardware-independent and commonly used as a billing metric for LLMs [VelascoTOG2025].
+We measure the number of generated tokens rather than runtime because token counts are hardware-independent and commonly used as a billing metric for LLMs [70].
 In contrast, runtime depends heavily on hardware factors such as GPU type, the number of GPUs used, their distribution across cluster nodes, and the intra- and inter-node interconnects.
 For each configuration, we report both the token count of the final valid schedule, which excludes retries, and the total tokens generated across all attempts, including failed ones.
-The table also indicates, for each junction architecture, the number of configurations that produced valid schedules and how many of these schedules required fewer shuttling operations than the classical compiler with the same configuration.
-In the following, we first present the results obtained by the LLMs for the linear and branched one-dimensional architectures with junctions, comparing them with each other and with the classical compiler.
+The table also indicates, for each junction architecture, the number of configurations that produced valid schedules and how many of these schedules required fewer shuttling operations than the baseline compiler with the same configuration.
+We now present the results obtained by the LLMs for the linear and branched one-dimensional architectures with junctions, comparing them with each other and with the baseline compilers [71, 30].
 Subsequently, in Sec.
@@
 The token count is given once without retries, where only the tokens of the final valid solution are included, and once with retries, where all tokens generated during invalid attempts are also counted.
-For the junction architecture, we additionally report how many configurations produced valid schedules and, among those, how many achieved fewer operations than the classical compiler on the same configuration.
+For the junction architecture, we additionally report how many configurations produced valid schedules and, among those, how many achieved fewer operations than the baseline compiler on the same configuration.
 Archi- tecture Parameter LLaMA 3.2-3B Qwen 3-4B Qwen 3-14B DeepSeek-7B Gemma 3-27B
@@
 4.2.1 Linear and branched one-dimensional architectures
-For 4mod5-bdd_287 on the linear architecture, DeepSeek, LLaMA at $T=0.7$, and Qwen-4B at $T=1.0$ produced schedules requiring fewer operations than the classical compiler, whereas Qwen-14B and Gemma consistently generated schedules with higher overhead.
-The best result was obtained by Qwen-4B at $T=1.0$, reducing the operation count by 17 %, indicating that the smaller Qwen model performs better than the larger one on this architecture.
-However, Qwen-4B also yielded the worst result at $T=0.7$, requiring 76 % more operations than the classical baseline.
+For 4mod5-bdd_287 on the linear architecture, DeepSeek, LLaMA at $T=0.7$, and Qwen-4B at $T=1.0$ produced schedules requiring fewer operations as compared with the baseline, whereas Qwen-14B and Gemma consistently generated schedules with higher overhead.
+The best result was obtained by Qwen-4B at $T=1.0$, reducing the operation effort by 15 %, indicating that the smaller Qwen model performs better than the larger one on this architecture.
+However, Qwen-4B also yielded the worst result at $T=0.7$, requiring 76 % more operations than the baseline.
 Across all models, 2 to 26 retries were needed, and valid schedules contained between 60,440 and 143,057 response tokens.
@@
 Notably, the best-performing Qwen-4B run also produced the smallest final token count, but its high number of 26 retries led to the largest overall increase when considering all generated tokens.
-For the junction architecture, LLaMA produced valid schedules for all ten tested configurations and, at $T=1.0$, outperformed the classical compiler in three cases.
-Qwen-14B produced valid schedules for all configurations except $s=1$, $d=4$ and surpassed the classical baseline for $s=5$, $d=1$ at both temperatures.
-Qwen-4B generated valid schedules for eight configurations and outperformed the classical compiler for the same configuration as Qwen-14B, but only at $T=1.0$.
-Gemma yielded valid schedules for seven configurations at $T=1.0$ and five at $T=0.7$, while DeepSeek produced valid schedules for only five configurations at $T=1.0$ and one at $T=0.7$; neither model surpassed the classical baseline.
-LLaMA’s best result, obtained for $s=2$, $d=3$ at $T=1.0$, required 4 % fewer operations than the best classical configuration $s=1$, $d=3$ and required only a single retry, leading to similar token counts with and without retries.
-In contrast, the best results of the other models required 7 % to 12 % more operations than the classical compiler and involved substantially more retries, ranging from 9 to 40.
+For the junction architecture, LLaMA produced valid schedules for all ten tested configurations and, at $T=1.0$, outperformed the baseline in three cases.
+Qwen-14B produced valid schedules for all configurations except $s=1$, $d=4$ and surpassed the baseline for $s=5$, $d=1$ at $T=0.7$ and $T=1.0$.
+Qwen-4B generated valid schedules for eight configurations and outperformed the baseline for the same configuration as Qwen-14B, but only at $T=1.0$.
+Gemma yielded valid schedules for seven configurations at $T=1.0$ and five at $T=0.7$, while DeepSeek produced valid schedules for only five configurations at $T=1.0$ and one at $T=0.7$; neither model surpassed the baseline.
+LLaMA’s best result, obtained for $s=2$, $d=3$ at $T=1.0$, required 4 % fewer operations than the best baseline configuration $s=1$, $d=3$ and needed only a single retry, leading to similar token counts with and without retries.
+In contrast, the best results of the other models required 7 % to 12 % more operations than the baseline and involved substantially more retries, ranging from 9 to 40.
 This resulted in token counts up to 5.1 times higher when retries were included.
@@
 However, Qwen-14B, DeepSeek, and Gemma failed to generate any complete schedules for the two larger circuits.
-For mini_alu_305 on the linear architecture, Qwen-4B at $T=1.0$ was the only model to produce a complete schedule, requiring 3,124 shuttling operations, corresponding to a 31 % increase over the classical compiler.
+For mini_alu_305 on the linear architecture, Qwen-4B at $T=1.0$ was the only model to produce a complete schedule, requiring 3,124 shuttling operations, corresponding to a 31 % increase over the baseline.
 The resulting schedule contained 555,430 tokens, and its computation required 33 retries, increasing the total number of generated tokens by a factor of 7.7.
@@
 Consequently, the models frequently reach trap configurations from which continuing schedule generation becomes infeasible.
-Moreover, since our compiler generates schedules gate by gate, the likelihood of encountering unrecoverable configurations grows with circuit depth.
+Moreover, since our compiler generates schedules gate after gate, the likelihood of encountering unrecoverable configurations grows with circuit depth.
 Nevertheless, the models were able to execute between 70 and 165 of the 263 gates, indicating that substantial partial progress remains achievable.
@@
 On the junction architecture, Gemma, Qwen-14B, and DeepSeek again failed to complete any schedule, executing at most 126, 85 and 69 gates, respectively.
-Qwen-4B produced a valid result only for the configuration $s=5$, $d=1$ at $T=1.0$, requiring 71 % more operations than the best classical configuration $s=2$, $d=1$ with 1,624 operations.
+Qwen-4B produced a valid result only for the configuration $s=5$, $d=1$ at $T=1.0$, requiring 71 % more operations than the best baseline configuration $s=2$, $d=1$ with 1,624 operations.
 This run required 58 retries, increasing the total number of generated tokens to 3.5 times the size of the final schedule.
 Among the failed runs, Qwen-4B at $T=0.7$ performed best, executing 168 gates before failure for the configuration $s=4$, $d=2$.
-However, LLaMA performed substantially better: of the 18 tested configurations, 15 at $T=0.7$ and 11 at $T=1.0$ yielded valid schedules, with one and two of these, respectively, outperforming the classical compiler.
+However, LLaMA performed substantially better: of the 18 tested configurations, 15 at $T=0.7$ and 11 at $T=1.0$ yielded valid schedules, with one and two of these, respectively, outperforming the baseline.
 Unlike the linear architecture, where LLaMA failed, the presence of stacks in the junction architecture provides intermediate storage locations that simplify reconfiguration and reduce the need for recursive operations.
-Relative to the best classical configuration, LLaMA achieved the best results for the configurations $s=2$, $d=3$ at $T=0.7$ and $s=2$, $d=2$ at $T=1.0$ requiring only 11 % more operations.
+Relative to the best baseline configuration, LLaMA achieved the best results for the configurations $s=2$, $d=3$ at $T=0.7$ and $s=2$, $d=2$ at $T=1.0$ requiring only 11 % more operations.
 These runs required 26 and 29 retries, respectively, roughly doubling the number of generated tokens compared to the final schedules.
@@
 DeepSeek failed after approximately 60 gates.
-Gemma’s performance strongly depended on the temperature: at $T=0.7$ it failed after executing 88 gates, while at $T=1.0$ it executed only 43 gates before failing.
+Gemma’s performance strongly depended on $T$: at $T=0.7$ it failed after executing 88 gates, while at $T=1.0$ it executed only 43 gates before failing.
 Notably, LLaMA’s performance indicates that, in principle, shuttling schedules for 16-qubit circuits are attainable, although robustness remains a limiting factor.
@@
 As for mini_alu_305, LLaMA also performed best here.
-Out of 29 tested configurations, LLaMA produced valid schedules for one at $T=0.7$ and two at $T=1.0$, but none outperformed the classical compiler.
-The best classical configuration $s=2$, $d=1$ required 1,716 operations, whereas LLaMA’s best result achieved with configuration $s=3$, $d=1$ at $T=1.0$ required 18 % more operations.
+Out of 29 tested configurations, LLaMA produced valid schedules for one at $T=0.7$ and two at $T=1.0$, but none outperformed the baseline.
+The best baseline configuration $s=2$, $d=1$ required 1,716 operations, whereas LLaMA’s best result achieved with configuration $s=3$, $d=1$ at $T=1.0$ required 18 % more operations.
 This run required 104 retries, increasing the number of generated tokens by a factor of 4.6 relative to the final schedule.
 4.2.2 Unseen architectures
-In addition to linear and branched one-dimensional architectures with junctions, we evaluated the generalization capabilities of the fine-tuned LLMs on the unseen architectures shown in Fig.
-5.
-While the layouts in 5(a) and 5(b) feature three-way junctions, similar to the branched architectures used during training, the architecture in 5(c) includes four-way junctions.
-For the architectures shown in 5(a) and 5(b), none of the models generated a valid shuttling schedule for any of the evaluated circuits.
+In addition to linear and branched one-dimensional architectures with junctions, we evaluated the generalization capabilities of the fine-tuned LLMs on the unseen architectures shown in Fig. 5.
+The layouts in 5(a) and 5(b) feature three-way junctions, similar to the branched architectures used during training.
+For all those architectures, none of the models generated a valid shuttling schedule for any of the evaluated circuits.
 While on the first architecture, all models failed immediately during schedule generation, on the second one, the models were able to execute up to four gates before failing.
-In contrast, one model succeeded in generating a valid schedule for the architecture shown in 5(c) for the circuit 4mod5-bdd_287.
+In contrast, one model successfully generated a valid schedule for the four-way junction architecture shown in 5(c) for the circuit 4mod5-bdd_287.
 LLaMA at $T=0.7$ produced a valid schedule with 569 shuttling operations but required 55 retries, increasing the total token count by a factor of 6.0 relative to the final schedule.
@@
 Nevertheless, partial progress was observed: Qwen-14B executed up to 52 gates, Qwen-4B up to 33 gates, and Gemma at $T=1.0$ 16 gates, while Gemma at $T=0.7$ and DeepSeek failed at an early stage.
-Consistent with the results on linear and branched architectures, performance on the four-way junction architecture degraded as the number of qubits increased.
+Consistent with the results before, performance on the four-way junction architecture degraded as the number of qubits increased.
 For the circuit mini_alu_305, no model generated a valid schedule.
@@
 These results further highlight that shuttling schedule generation becomes increasingly challenging as the number of qubits grows.
-Overall, our results show that fine-tuned LLMs can generate valid shuttling schedules and can, in some cases, outperform classical compilers.
-However, producing complete schedules becomes increasingly challenging as the number of qubits and gates increases.
-The models frequently enter invalid states, requiring retries that substantially inflate token counts without contributing to the final output.
+Overall, our results show that fine-tuned LLMs can generate valid shuttling schedules and can, in some cases, outperform standard heuristics.
+However, producing complete schedules becomes increasingly challenging with the width and depth of the algorithm.
+Models may enter invalid states, requiring retries that substantially inflate token counts without contributing to the final output.
 Generalization to previously unseen architectures remains limited.
-Across all experiments, LLaMA demonstrated the most stable performance, whereas Qwen, DeepSeek, and Gemma failed more frequently and at earlier stages as circuit size increased.
+Across all experiments, LLaMA demonstrated the most stable performance, whereas Qwen, DeepSeek, and Gemma failed more frequent and at earlier stages as circuit size increased.
 When comparing the both Qwen models, it can be noticed that the smaller model with 4 billion parameters produced better results than the the larger model with 14 billion parameters.
@@
 We fine-tuned several LLMs on datasets derived from shuttling schedules generated by classical compilers for linear and branched one-dimensional trap architectures and evaluated them on circuits of varying sizes.
-For small circuits on both linear and junction layouts, the models were able to generate valid schedules and in some cases even required fewer operations than the classical compilers.
-As circuit size increased, generating complete shuttling schedules became substantially more difficult.
-Larger numbers of qubits make trap reconfiguration increasingly complex and often require long, recursive sequences of shuttling operations.
-As a result, the models frequently succeeded in generating valid operations only for an initial portion of the circuit before reaching trap configurations from which continuation would require operations beyond their generation capabilities.
-Since our compiler constructs schedules gate by gate, the likelihood of entering such unrecoverable configurations grows with circuit depth.
-These results indicate that LLM-based shuttling compilation is feasible in principle but remains limited in robustness.
-Evaluations on previously unseen trap layouts further revealed a sharp performance degradation, indicating that the models generalize poorly beyond the architectures encountered during training.
-To further improve the capabilities of our compiler, we plan to enhance the training pipeline in two stages: first using Direct Preference Optimization (DPO) [RafailovSMEMF23], and later using Gradient Regularized Policy Optimization (GRPO) [OpenAI24].
+For small circuits on both linear and junction layouts, the models were able to generate valid schedules and in some cases even required fewer operations as compared to the standard heuristics [71, 30].
+To improve the capabilities of LLM-compilers, we plan to enhance the training pipeline in two stages: first using Direct Preference Optimization (DPO) [50], and later using Gradient Regularized Policy Optimization (GRPO) [43].
 Applying DPO requires generating multiple candidate shuttling operations at each step so that the model can learn which alternatives are preferable.
@@
 While the runtime of this approach grows exponential, it provides valuable multi-step supervision.
-Building on this, GRPO offers a framework for reinforcement learning with human feedback (RLHF) [ChristianoLBMSD17].
+Building on this, GRPO offers a framework for reinforcement learning with human feedback (RLHF) [8].
 This would require a continuous reward function that penalizes invalid schedules and rewards valid schedules with fewer operations, again ideally using a lookahead-based evaluation.
 Due to the large number of shuttling generations needed for DPO and the additional computational overhead of GRPO, scaling these methods will require access to more GPUs than the eight H100 GPUs available for the present work.
-Overall, our work establishes the first foundation for LLM-based shuttling compilers for trapped-ion quantum computers and identifies promising directions to significantly extend their performance and reliability in future research.
+Overall, our work establishes the first LLM-based shuttling compilers for trapped-ion quantum computers and may paving a way to extend performance and reliability.
 Acknowledgments
-We acknowledge funding by the German Federal Ministry of Education and Research (BMBF) via the VDI within the project IQuAn.
+We acknowledge funding by the German Federal Ministry of Research, Technology and Space (BMFTR) within the project IQuAn, ATIQ and SYNQ, and the DFG SPP2514.
 We thank the data center (ZDV) of the Johannes Gutenberg University for the use of their Mogon NHR and Mogon KI clusters for evaluating our LLMs.